### v1

# Multi-scale Restoration of Missing Data in Optical Time-series Images with Masked Spatial-Temporal Attention Network

Abstract
Due to factors such as thick cloud cover and sensor limitations, remote sensing images often suffer from significant missing data, resulting in incomplete time-series information. Existing methods for imputing missing values in remote sensing images do not fully exploit spatio-temporal auxiliary information, leading to limited accuracy in restoration. Therefore, this paper proposes a novel deep learning-based approach called MS2TAN (Multi-scale Masked Spatial-Temporal Attention Network), for reconstructing time-series remote sensing images. Firstly, we introduce an efficient spatio-temporal feature extractor based on Masked Spatial-Temporal Attention (MSTA), to obtain high-quality representations of the spatio-temporal neighborhood features in the missing regions. Secondly, a Multi-scale Restoration Network consisting of the MSTA-based Feature Extractors, is employed to progressively refine the missing values by exploring spatio-temporal neighborhood features at different scales. Thirdly, we propose a “Pixel-Structure-Perception” Multi-Objective Joint Optimization method to enhance the visual effects of the reconstruction results from multiple perspectives and preserve more texture structures. Furthermore, the proposed method reconstructs missing values in all input temporal phases in parallel (i.e., Multi-In Multi-Out), achieving higher processing efficiency. Finally, experimental evaluations on two typical missing data restoration tasks across multiple research areas demonstrate that the proposed method outperforms state-of-the-art methods with an improvement of 0.40dB/1.17dB in mean peak signal-to-noise ratio (mPSNR) and 3.77‰/9.41‰ in mean structural similarity (mSSIM), while exhibiting stronger texture and structural consistency.
Index Terms:
missing data restoration, time-series remote sensing images, masked spatial-temporal attention, multi-scale restoration, multi-objective joint optimization.
I Introduction
Over the past few decades, remote sensing data has been extensively used in various industries. Among them, high spatial resolution remote sensing imagery is particularly beneficial for applications such as vegetation monitoring, land cover mapping, and land cover change detection. However, the fine spatial resolution images suffer from inevitable information loss caused by internal factors (e.g., sensor malfunction) and external factors (e.g., atmospheric conditions), which restrict their applications in different domains [1].
This paper addresses the issue of missing data in remote sensing image involving multiple spectra. Common tasks include resolving the Landsat-7 ETM+ sensor scan line corrector off (SLC-off) problem and removing thick clouds. The key challenge is to estimate the missing regions and fill the gaps with predicted pixels, ensuring visual and semantic consistency with the surrounding pixels to enhance data usability.
Researchers have proposed various methods to recover missing data in remote sensing images. Early approaches to missing value restoration can be broadly categorized into three types [1]: spatial-based, temporal-based, and spatio-temporal-based data recovery. These methods have shown promising results in specific scenarios with low resolution and low missing rates. However, most of them rely on linear models and struggle to handle complex and detailed scenes. Moreover, due to limited reference information, the generated images often exhibit blurriness and lack continuous textures.
In recent years, the rapid advancement of deep learning theory and computer hardware [2] has led to significant progress in remote sensing image restoration using deep learning-based methods. These methods can be broadly categorized into two types: CNN-based [3] and Transformer-based [4] approaches. Compared to traditional statistical models, CNNs exhibit strong non-linear expressive power, allowing for efficient extraction of spatial features from remote sensing data and significantly improving the accuracy of image restoration. However, these structures struggle to fully exploit time-series information, resulting in a bottleneck in restoration accuracy. Transformers, based on self-attention mechanisms, possess a global receptive field, enabling comprehensive and efficient utilization of both temporal sequences and images [5] to enhance reconstruction results [6]. However, due to the high resolution and long time-series of remote sensing images, token sequences become excessively long, leading to extremely high complexity in attention computations.
To efficiently mine spatio-temporal information in remote sensing images, we apply self-attention mechanism separately in the temporal and spatial dimensions and alternate between the two [7], greatly reducing computational complexity. To address the problem of significant distributional differences between missing and non-missing values in remote sensing images, we apply missing values mask and diagonal mask to the attention matrix [8], proposing Masked Spatial-Temporal Attention (MSTA) to enhance the expressive power of spatio-temporal attention and optimize the spectral discrepancy at the transition regions.
For full exploitation of the spatio-temporal neighborhood features at different scales [9], we further propose a Multi-scale Restoration Network. The network consists of MSTA-based Feature Extractors with different embedding scales, which progressively refine the reconstruction of missing information from coarse to fine granularity levels, achieving higher restoration accuracy [10]. For model training, we propose a “Pixel-Structure-Perception” Multi-Objective Joint Optimization method, using pixel-wise loss as the basic loss for the restoration task and incorporating structural loss [11] and perceptual loss [12] to optimize the model’s results from the perspectives of structure, texture, shape, and spatial relations, thereby achieving high-quality image inpainting.
Finally, we conducted simulated and real data experiments to validate the performance of the proposed method for the ETM+ SLC-off and thick cloud removal tasks. The results from six research areas indicate that our method outperforms competitors. In addition, a validation study demonstrates the effectiveness of the core components.
In summary, our main contributions are as follows:
  1. 1.
    We propose a deep learning-based approach called MS2TAN, for repairing missing information in time-series remote sensing images. Our method utilizes a Multi-scale Restoration Network with MSTA-based Feature Extractors to learn the non-linear mapping between missing and complete data. Both simulated and real data experiments show that proposed MS2TAN outperforms state-of-the-art methods across multiple research areas.
  2. 2.
    To address the challenges of high resolution and long temporal sequences in remote sensing data, we introduce the Masked Spatial-Temporal Attention (MSTA) mechanism. The MSTA efficiently extracts spatio-temporal features, improves the utilization of spatio-temporal contextual information, reduces color transition artifacts at the boundaries of missing value regions, and significantly reduces the time complexity of self-attention.
  3. 3.
    For model training, we propose a “Pixel-Structure-Perception” Multi-Objective Joint Optimization method. This method considers pixel-wise reconstruction error, structural reconstruction error, and perceptual error, resulting in restoration results with enhanced visual quality and preserved texture and structural details.
The rest of this paper is organized as follows. In Section II, we review the existing methods for reconstructing missing information in remote sensing images. The network architecture and methodology details of our proposed model are presented in Section III. In Section IV, we showcase the results of missing data reconstruction in both simulated and real data experiments, compare them with mainstream methods, and validate the effectiveness of each component through validation study. Finally, our conclusions and future prospects are summarized in Section V.
II Related Works
II-A Traditional methods
Early research on the restoration of missing data in remote sensing images can be roughly divided into three categories: spatial-based, temporal-based, and spatio-temporal hybrid methods.
II-A1 Spatial-based Methods
Spatial-based methods rely solely on the valid information within the image itself to predict the missing data. The most commonly used approach is spatial interpolation methods [13]. Additionally, methods based on partial differential equations (PDE) [14] and variational methods [15, 16, 17] have also been utilized for reconstructing missing values. Furthermore, patch-based methods have found extensive application [18, 19, 20]. In general, spatial-based methods are suitable for reconstructing small missing areas or regions with regular textures. However, the accuracy of the reconstruction cannot be guaranteed, particularly for large regions or complex textures.
II-A2 Temporal-based Methods
Temporal-based methods utilize observations of the same location at different times from satellites to restore missing data. These methods include histogram matching-based approaches [21], temporal interpolation-based methods [22], replacement-based methods [23], and regression-based methods [24, 25, 26], among others. However, the differences between different time phases restrict the application of these methods.
II-A3 Spatio-temporal-based Methods
To overcome these limitations, spatio-temporal-based methods integrate the spatial and temporal correlations to reconstruct missing data under various conditions. For instance, improved nearest neighbor pixel interpolation methods [27], methods based on spatio-temporal Markov random field model [28], spatio-temporal weighted regression model [29], methods based on group sparse representation [30], and methods based on low-rank tensor decomposition [31, 32] have been proposed. However, most of these methods rely on linear models and struggle to handle complex and intricate scenes.
II-B Deep learning-based methods
In recent years, deep learning methods based on CNN and Transformer have been widely applied.
II-B1 CNN-based methods
CNNs have shown high efficiency in extracting spatial features from remote sensing data, leading to significant improvements in the accuracy of remote sensing image restoration. Malek et al. [33] applied a context encoder [34] to reconstruct thick clouds in remote sensing images. CNNs combined with GAN structures [35, 36] were used to generate realistic reconstructed images. Zhang et al. [37] proposed a CNN-based Spatio-Temporal Spectra (STS-CNN) framework, which was further developed into a progressive spatio-temporal patch grouping framework [38]. CNNs that incorporate temporal inputs through channel-wise concatenation [39] merge feature maps of target images and temporal images, introducing auxiliary information for missing data restoration. However, CNNs lack a true understanding of time series, which hinders the efficient integration of temporal and spatial information in these methods and limits the utilization of long temporal sequences as auxiliary information, ultimately affecting the restoration accuracy.
II-B2 Transformer-based methods
Transformer [4] employs self-attention mechanism to process text sequences and capture global contextual information. It has been proven that Transformer-based methods, such as Visual Transformer (ViT) [5], also exhibit powerful performance in the field of image processing. The sequence modeling capability of self-attention mechanism can be leveraged to assist in high-precision restoration of missing values in remote sensing images. Xu et al. [40] applied spatial self-attention to feature maps to capture the distribution of cloud thickness. Christopoulos et al. [41] utilized axial attention to remove thick clouds in remote sensing images. Recently, Liu et al. [42] used spatial attention and channel attention to remove cloud cover in the images. Li et al. [43] introduced SAR auxiliary data and synthesized cloud-free images using a Transformer-based GAN.
However, the spatial attention used in ViT requires computing the correlations between all pairs of patches. In the context of processing time-series remote sensing images, the images often have high resolutions and long time-series, resulting in a large number of tokens and extremely high computational complexity. Bertasius et al. [7] conducted a detailed comparison of various forms of spatial-temporal attention and proposed separated spatial-temporal attention that achieves a balance between efficiency and performance. We further improved the separated spatial-temporal attention by introducing missing value masks and diagonal masks [8], resulting in Masked Spatial-Temporal Attention (MSTA), which exhibits superior performance in the task of missing value restoration. Compared to CNN-based methods and original ViT methods, MSTA enables efficient processing of spatio-temporal information and leverages long-term temporal information to assist in missing value reconstruction, leading to more precise reconstruction results.
Figure 1: The overall flowchart of the proposed method consists of two main components: a Multi-scale Restoration Network (named MS2TAN) with MSTA-based Feature Extractors, and a “Pixel-Structure-Perception” Multi-Objective Joint Optimization method.
III Methodology
III-A Problem Definition and Overall Framework
III-A1 Problem Definition
The original time-series remote sensing image with missing values can be represented as $X\in\mathbb{R}^{T\times C\times H\times W}$, where $T$ denotes the length of the time series, $C$ represents the number of channels, and $H$ and $W$ denote the spatial dimensions of the region. $X_{(t,c,i,j)}$ denotes the value of channel $c$ at position $(i,j)$ at time $t$. To account for the missing values, we introduce the missing value hint tensor $M\in\mathbb{R}^{T\times C\times H\times W}$, which is defined as shown in Equation (1). In practical applications, the missing value hint tensor $M$ can be obtained from QA band or cloud detection algorithms such as Fmask [44].
$M_{(t,c,i,j)}=\left\{\begin{matrix}0&\text{if}~{}X_{(t,c,i,j)}~{}\text{is~{}missing},\\ 1&\text{if}~{}X_{(t,c,i,j)}~{}\text{is~{}observed}.\end{matrix}\right.$ (1)
The target sequence $Y$ in $\mathbb{R}^{T\times C\times H\times W}$ represents the actual data, while $\widetilde{Y}$ in $\mathbb{R}^{T\times C\times H\times W}$ represents the data repaired by the model. Therefore, the definition of the multi-temporal remote sensing image reconstruction model is given by equation (2). Here, the output $\widetilde{Y}$ of the model is the initial reconstruction result, and $\odot$ denotes the Hadamard product. By replacing the observed real values in $X$ with $\widetilde{Y}$, we obtain the final reconstruction result $\widetilde{Y}_{\mathrm{out}}$ as shown in equation (3).
$\widetilde{Y}=\operatorname{Model}(X,M)$ (2)
$\widetilde{Y}_{\mathrm{out}}=\widetilde{Y}\odot(1-M)+X\odot M$ (3)
Our objective is to make $\widetilde{Y}$ closely match the data distribution of $Y$ in order to achieve high-quality reconstructed images.
III-A2 Overall Framework
The proposed framework for time-series image recovery consists of a Multi-scale Restoration Network (called MS2TAN) with MSTA-based Feature Extractors and a “Pixel-Structure-Perception” Multi-Objective Joint Optimization method, as depicted in Fig. 1.
The framework utilizes the MS2TAN to learn the non-linear mapping from non-missing information to missing information. It employs the MSTA-based Feature Extractor (MFE) at different scales to extract temporal and spatial features for reconstruction. Finally, the network parameters are optimized using the “Pixel-Structure-Perception” Multi-Objective Joint Optimization method, and the trained parameters are used for inference. The details of these components will be discussed in the following sections III-B, III-C and III-D.
Figure 2: Illustration of the structure of the MSTA-based Feature Extractor (MFE). The input sequence is augmented with positional encoding to incorporate positional information. It then undergoes $L$ layers of Masked Spatial-Temporal Attention operations to obtain the output sequence, which combines temporal and spatial contextual features. The Masked Temporal/Spatial Attention, introduced below the image, respectively employ masked self-attention to capture the data distribution features in the Temporal/Spatial neighborhood.
III-B MSTA-based Feature Extractor
The structure of the MSTA-based Feature Extractor (MFE) module is illustrated in Fig. 2. It consists of position encoding and a cascade of $L$ Masked Spatial-Temporal Attention (MSTA) units. The input to this module is a high-dimensional feature vector $E\in\mathbb{R}^{TN\times d_{\mathrm{emb}}}$, which is obtained by embedding the time-series remote sensing image $X$ and the missing information mask $M$ into blocks (as described in detail in Section III-C1). Here, $T$ represents the time-series length, $N$ represents the number of patches in a single remote sensing image, and $d_{\mathrm{emb}}$ represents the dimension of the token vector corresponding to each patch. Subsequently, the position encoding introduces spatial-temporal positional semantics, followed by $L$ layers of MSTA units, and ultimately produces a sequence with incorporated spatio-temporal correlation features.
III-B1 Masked Self-Attention
The Transformer model utilizes self-attention mechanism to model sequences. It maps the input to query vector Q, key vector K, and value vector V. The attention score between Q and K is computed using Scaled Dot-Product, followed by the application of the Softmax function to obtain attention weights A. The final output is the attention-weighted V, as shown in Equation 4.
$\begin{gathered}\mathrm{H}=\mathrm{A}\mathrm{V}=\mathrm{Softmax}\left(\frac{\mathrm{Q}\cdot\mathrm{K}^{\top}}{\sqrt{d_{k}}}\right)\cdot\mathrm{V}\\ \mathrm{where}\ \mathrm{[Q,K,V]}=X\cdot W_{\mathrm{QKV}}.\end{gathered}$ (4)
To address the challenge of disparate distribution of missing values in remote sensing data and enhance feature extraction capability, we introduce missing value masks and diagonal masks into self-attention.
  • The missing value mask sets the attention scores from patches with too high missing rate to other patches as $\infty$- ∞, effectively masking the influence of missing values. This enables the model to focus on extracting useful information from non-missing data to fill in the gaps, without being affected by the missing data.
  • The diagonal mask sets the diagonal of the attention matrix as $\infty$- ∞, preventing each step from attending to itself and forcing it to rely on the other $TN-1$ steps for inference. This helps capture the spatio-temporal feature correlation in high-dimensional space.
More specifically, for the attention score matrix $\mathrm{Sc}\in\mathbb{R}^{TN\times TN}$, the attention mask operation $\mathrm{ApplyMask}$ is defined as shown in Equation (5).
$\begin{gathered}\left.\mathrm{ApplyMask}\left(\mathrm{Sc}\right)_{\left(i,j\right)}=\left\{\begin{array}[]{ll}-\infty&\mathrm{Mask}(i,j)=1,\\ \mathrm{Sc}_{\left(i,j\right)}&\mathrm{Mask}(i,j)\neq 1.\end{array}\right.\right.\\ \mathrm{Mask}(i,j)=1\ \textbf{only if}\ (\mathrm{MR}(E(i))>C_{\mathrm{max}}\ \textbf{or}\ i=j)\end{gathered}$ (5)
Here, $Mask()$ checks whether the mask condition is satisfied, $\mathrm{MR}(e)$ represents the missing rate of the corresponding patch $e$ (which can be calculated from the hint tensor $M$), and $C_{\mathrm{max}}$ is a hyperparameter that controls the maximum allowable missing rate. Applying Equation (5) to the attention scores in Equation (4), we obtain the expression for a single masked attention head, as shown in Equation (6). Finally, by linearly projecting and combining multiple attention heads with the residual connection, we obtain the output of the masked self-attention, as shown in Equation (7).
$\mathrm{H}^{(i)}=\mathrm{A}^{\prime}\mathrm{V}=\mathrm{Softmax}\left(\mathrm{ApplyMask}\left(\frac{\mathrm{Q}\cdot\mathrm{K}^{\top}}{\sqrt{d_{k}}}\right)\right)\cdot\mathrm{V}$ (6)
$\mathrm{MaskedAttn}(X,M)=\mathrm{Proj}\left(\left[\mathrm{H}^{(1)},\mathrm{H}^{(2)},\cdots,\mathrm{H}^{(h)}\right]\right)+X$ (7)
In Equation (7), $\mathrm{A}^{\prime}$ denotes the attention weights post mask application, $\mathrm{MaskedAttn}()$ represents the masked self-attention operation, $\mathrm{Proj}$ stands for projection head, and $h$ denotes the number of attention heads.
Figure 3: Schematic of MSTA. Taking the blue patch as an example, patches at the same spatial position are considered as temporal auxiliaries, while patches at the same time are regarded as spatial auxiliaries. Additionally, patches with excessively high missing rates will be excluded from the receptive field to avoid the influence of distributional differences on model learning.
III-B2 Masked Spatial-Temporal Attention
Masked Spatial-Temporal Attention (MSTA) captures the spatio-temporal correlations within the input sequence, with its masking mechanism and receptive field depicted in Fig. 3. Specifically, MSTA comprises Masked Temporal Attention (MTA), Masked Spatial Attention (MSA), Layer Normalization (LN), and Feed-Forward Network (FFN). These components are elaborated upon below.
For an input sequence $E\in\mathbb{R}^{TN\times d_{\mathrm{emb}}}$, MTA first reshapes $E$ to place the time dimension and the feature dimension at the end, obtaining a sequence $e_{t}$ of length $T$ and $d_{\mathrm{emb}}$ dimensions. Subsequently, for each position in space, MaskedAttn operation is performed along the temporal direction to obtain the output of MTA, as shown in Equation (8).
$\begin{gathered}e_{t}=\mathrm{Reshape}(E)\in\mathbb{R}^{N\times T\times d_{\mathrm{emb}}}\\ \mathrm{MTA}(E,M)=\mathrm{Reshape}\left(\mathrm{MaskedAttn}\left(e_{t},M\right)\right)\end{gathered}$ (8)
Similarly, after the corresponding dimension transformation of the input sequence, MaskedAttn operation is conducted along the spatial direction for each temporal image, yielding the output of MSA, as shown in Equation (9).
$\begin{gathered}e_{s}=\mathrm{Reshape}(E)\in\mathbb{R}^{T\times N\times d_{\mathrm{emb}}}\\ \mathrm{MSA}(E,M)=\mathrm{Reshape}\left(\mathrm{MaskedAttn}\left(e_{s},M\right)\right)\end{gathered}$ (9)
FFN consists of two linear layers separated by a ReLU activation function, with a residual connection established between the input and output. The expression for FFN is presented in Equation (10).
$\mathrm{FFN}\left(X\right)=\mathrm{Linear}\left(\mathrm{ReLU}\left(\mathrm{Linear}\left(X\right)\right)\right)+X$ (10)
MSTA sequentially applies MTA and MSA to the input token sequence, exploring the spatio-temporal correlations among patches. Subsequently, FFN is used to fuse spatio-temporal features and introduce non-linear transformations, as shown in Equation (11).
$\begin{gathered}U=\operatorname{MTA}(\operatorname{LN}(E),M)\\ V=\operatorname{MSA}(\operatorname{LN}(U),M)\\ \operatorname{MSTA}(E,M)=\operatorname{FFN}\left(\operatorname{LN}\left(V\right)\right)\end{gathered}$ (11)
Figure 4: The network structure diagram of proposed MS2TAN (Multi-Scale Masked Spatial-Temporal Attention Network), which includes components such as Embedding, MFE, Unembedding, and Observed Value Replacement.
III-B3 Positional Encoding
In the Transformer, positional encoding is added to the input sequence to introduce spatio-temporal positional semantics, as shown in Equation (12), where $\operatorname{PosEnc}(pos,dim)$ represents the value of the $dim$-th dimension of the positional encoding for the $pos$-th element in the sequence.
$\begin{gathered}\operatorname{PosEnc}(pos,2i)=\sin\left({pos}\cdot{10000^{-{2i}/{d_{\mathrm{emb}}}}}\right)\\ \quad\operatorname{PosEnc}(pos,2i+1)=\cos\left({pos}\cdot{10000^{-{2i}/{d_{\mathrm{emb}}}}}\right)\end{gathered}$ (12)
After positional encoding, the input sequence will undergo $L$ layers of MSTA to explore deeper spatio-temporal features. Finally, the overall expression of the MFE module is shown in Equation (13), where the symbol L denotes stacking $L$ layers, and $\mathrm{MFE}()$ represents the MSTA-based Feature Extractor.
$\begin{gathered}\mathrm{MFE}\left(E,M\right)=\{\mathrm{MSTA}\left(E+\mathrm{PosEnc},M\right)\}^{L}\end{gathered}$ (13)
III-C Multi-scale Restoration Network
The MS2TAN (Multi-Scale Masked Spatial-Temporal Attention Network) is a multi-scale restoration network composed of $S$ residual-connected restoration modules and observed value replacement, as illustrated in Fig. 4. Each module consists of Patch Embedding, MFE, and Patch Unembedding. These restoration modules can capture the spatio-temporal correlations of pixels at different scales to predict missing values and ultimately replace observed values as outputs.
Figure 5: Illustration of Embedding and Unembedding processes. Embedding represents a series of transformations from left to right, and Unembedding represents the reverse.
III-C1 Embedding and Unembedding
The input to MS2TAN consists of two components: the image sequence $X_{\mathrm{in}}$ and the missing value hint tensor $M$. To incorporate the distributional properties of the missing values into the model learning process, $X_{\mathrm{in}}$ and $M$ are concatenated along the channel dimension to yield the merged input $X_{c}=\mathrm{Concat}(X_{\mathrm{in}},M)$, where $X_{c}=\left[x_{c}^{1},x_{c}^{2},\cdots,x_{c}^{T}\right]\in\mathbb{R}^{T\times C^{\prime}\times H\times W}$, where $C^{\prime}=C+1$ is the total number of channels after concatenation.
Following the approach of Vision Transformer (ViT), each image is partitioned into a series of patches, with the size of each patch determined by the hyperparameter $P$. For each temporal image, $N=HW/P^{2}$ patches are obtained. These patches are then aggregated and reorganized to form a patch sequence of length $TN$, with each patch represented as a $C^{\prime}P^{2}$-dimensional vector. Equation (14) delineates the processing of the $t$-th temporal image $x_{c}^{t}$.
$\begin{gathered}A^{t}=\left[p_{1}^{t},p_{2}^{t},\cdots,p_{N}^{t}\right]=\operatorname{Patchify}(x_{c}^{t},P)\\ \text{where}\ x_{c}^{t}\in\mathbb{R}^{C^{\prime}\times H\times W}\ \text{and}\ A^{t}\in\mathbb{R}^{N\times C^{\prime}P^{2}}\end{gathered}$ (14)
Next, a linear layer is utilized to project the vectors corresponding to each patch into a high-dimensional representation space. The images from different time steps are then aggregated to obtain the complete token sequence $E$, as depicted in Equation (15).
$\begin{gathered}E^{t}=\mathrm{Linear}(A^{t})\in\mathbb{R}^{N\times d_{\mathrm{emb}}}\\ E=\left[E^{1},E^{2},\cdots,E^{T}\right]\in\mathbb{R}^{TN\times d_{\mathrm{emb}}}\end{gathered}$ (15)
The process of transforming $X_{\mathrm{in}}$ and $M$ step by step to obtain $E$ as described above is referred to as Embedding, which can be represented by Equation (16). Conversely, if this operation is reversed, i.e., $E$ is restored to a patch sequence and then rearranged into an image, it is referred to as Unembedding, as shown in Equation (17). Fig. 5 illustrates the processes of Embedding and Unembedding on images.
$E=\mathrm{Embedding}\left(\mathrm{Concat}\left(X_{\mathrm{in}},M\right),P\right)$ (16)
$X_{\mathrm{out}}=\mathrm{Unembedding}\left(E,P\right)$ (17)
III-C2 MS2TAN
MS2TAN consists of components such as Embedding, MFE, and Unembedding, where the block size $P$ is an important hyperparameter, and the embedding scale of the $i$-th layer is denoted as $P^{\left(i\right)}$. The process of obtaining the embedding vector $\alpha$ from the input is described in Equation (18). Subsequently, MFE is utilized to mine deep spatio-temporal correlation features $\beta$, as shown in Equation (19). Finally, the features are decoded, unembedded back into image form, and added to the input, resulting in an intermediate result as depicted in Equation (20), where $\widetilde{Y}^{(i)}$ represents the $i$-th intermediate result.
$\alpha=\mathrm{Embedding}\left(\mathrm{Concat}\left(\widetilde{Y}^{(i-1)},M\right),P^{\left(i\right)}\right)$ (18)
$\beta=\mathrm{MFE}\left(\alpha,M\right)$ (19)
$\widetilde{Y}^{(i)}=\mathrm{Unembedding}\left(\beta,P^{\left(i\right)}\right)+\widetilde{Y}^{(i-1)}$ (20)
Let $\widetilde{Y}^{(0)}=X_{\mathrm{in}}$, and sequentially obtain $S$ intermediate results of the reconstruction. Eventually, the $S$-th result is taken as the output of MS2TAN, as shown in Equation (21).
[…]
III-C3 Observed Value Replacement
For the original reconstruction results of this network, denoted as $\widetilde{Y}$, the observed values are replaced with their non-missing parts to obtain the final result $\widetilde{Y}_{\mathrm{out}}$, as shown in Equation (22).
$\widetilde{Y}_{\mathrm{out}}=\widetilde{Y}\odot(1-M)+X\odot M$ (22)
III-D Multi-Objective Joint Optimization
To fully exploit the reconstruction capability of the model, this paper proposes a multi-objective joint optimization method to train the network parameters. This method utilizes the “Pixel-Structure-Perception” Multi-Objective Loss Function to optimize the results generated by the model from the perspectives of structure, color, texture, shape, and spatial relationships, achieving high-quality image inpainting.
III-D1 Pixel-wise Loss
Pixel-wise loss disregards the overall integrity of the image, treating the image as a collection of pixels, and comparing pixel by pixel to generate the image against the target image. It serves as the foundational loss for remote sensing image reconstruction tasks. Specifically, Equation (23) illustrates its calculation process.
${\mathcal{L}}_{\textrm{pixel-wise}}=\frac{1}{TCHW}\left|\left|\eta-y\right|\right|_{2}^{2}$ (23)
The pixel-wise loss calculated here includes both missing and observed parts. The reduction of the loss in the observed part does not directly improve the performance of the model. However, it does smooth the model output, making the model training process more stable, and hence is also included.
III-D2 Structural Loss
Pixel-wise reconstruction loss is commonly employed in various time-series and visual tasks but fails to consider the correlation between pixels. Structural reconstruction loss, based on the principle of structural similarity [45], measures the difference with the target image using the structural similarity index (SSIM), optimizing the visual consistency of the reconstruction results from the perspectives of structure, contrast, and luminance.
${\mathcal{L}}_{\mathrm{structural}}$ $=1-\mathrm{SSIM}\left(\eta,y\right)$ (24)
$=1-\frac{\left(2{\mu_{\eta}}{\mu_{y}}+C_{1}\right)\left(2\sigma_{\eta y}+C_{2}\right)}{\left(\mu_{\eta}^{2}+\mu_{y}^{2}+C_{1}\right)\left(\sigma_{\eta}^{2}+\sigma_{y}^{2}+C_{2}\right)}$
Equation (24) specifies the computation process of structural loss, where $\mu_{\eta}$ and $\mu_{y}$ denote the means of $\eta$ and $y$ respectively, $\sigma_{\eta}$ and $\sigma_{y}$ denote the variances of $\eta$ and $y$ respectively, $\sigma_{\eta y}$ denotes the covariance between $\eta$ and $y$, and $C_{1}$ and $C_{2}$ are constants.
III-D3 Perceptual Loss
Perceptual loss utilizes a pre-trained VGG16 network [46], as the feature extraction network to obtain perceptual feature vectors, followed by applying L2 loss to the feature vectors. The calculation of perceptual loss is shown in Equation (25), where $\psi\left(\right)$ represents the feature extraction network, and $d_{f}$ denotes the feature dimension.
${\mathcal{L}}_{\mathrm{perceptual}}=\frac{1}{d_{f}}\left|\left|\psi\left(\eta\right)-\psi\left(y\right)\right|\right|_{2}^{2}$ (25)
III-D4 Multi-Objective Loss Function
For each intermediate output $\widetilde{Y}^{(i)}$ of MS2TAN, the expression of the multi-objective loss function is shown in Equation (26), where $\lambda_{\mathrm{s}}$ and $\lambda_{\mathrm{p}}$ denote the weights of the structural loss and perceptual loss respectively. The final loss function considers the intermediate results of each layer, as shown in Equation (27).
$\mathcal{L}^{\left(i\right)}\left(\eta,y\right)=\mathcal{L}_{\text{pixel-wise}}+{\lambda_{\mathrm{s}}\mathcal{L}}_{\mathrm{struct}}+{\lambda_{\mathrm{p}}\mathcal{L}}_{\mathrm{perceptual}}$ (26)
$\mathcal{L}=\frac{1}{S}\sum_{i=1}^{m}\mathcal{L}^{\left(i\right)}\left(\widetilde{Y}^{(i)},Y\right)$ (27)
Figure 6: The locations of the research areas used to evaluate the proposed methodology are shown with examples. S1 to S6 correspond to the locations of the six research areas in Table I: Wuhan, Hengyang, Shenzhen, Dallas, Kern, and Urumqi.
TABLE I: Details of Six Research Areas
  Satellite Task Location Date Range Area Size and Resolution
  S1 Landsat-5 TM ETM+ SLC-off Wuhan, Hubei, China 2001-03 to 2003-10 (1720, 2040) , 30-m
S2 Landsat-5 TM ETM+ SLC-off Hengyang, Hunan, China 2003-04 to 2005-05 (1200, 1200) , 30-m
S3 Landsat-5 TM Thick Cloud Removal Shenzhen, Guangdong, China 2003-10 to 2004-11 (1680, 1680) , 30-m
S4 Landsat-8 OLI Thick Cloud Removal Dallas, Texas, United States 2014-01 to 2014-11 (2040, 2040) , 30-m
  S5 Landsat-7 ETM+ Real ETM+ SLC-off Kern, California, United States 2011-09 to 2011-11 (1720, 2040) , 30-m
S6 Landsat-5 TM Real Thick Cloud Removal Urumqi, Xinjiang, China 2006-06 to 2009-08 (1440, 1440) , 30-m
IV Experiment
IV-A Experiment Settings
To assess the performance of the proposed algorithm, we conducted experiments on both simulated and real-world datasets. Additionally, we investigated the effectiveness of each core component of the proposed method.
IV-A1 Datasets
Different datasets were tailored for different missing data reconstruction tasks. For instance, for simulated ETM+ SLC-off and thick cloud removal tasks, we selected two research areas each for model training and testing. For real-world SLC-off and thick cloud removal, we chose one research area each and evaluated the trained model within them. All research areas comprised multiple images captured at different times. We employed a sliding window approach with a stride of 60 and a size of 120$\times$120 to augment the sample amount, and then split them into training and testing sets. Specifically, Fig. 6 illustrates the spatial distribution and examples of each research area, while Table I provides detailed information on their data sources, tasks, time ranges, etc.
IV-A2 Comparison Algorithms and Evaluation Metrics
For the ETM+ SLC-off and thick cloud cover problems, we compared our method with LLHM [21], WLR [24], and STS-CNN [37]. These algorithms take a target image and a single temporal auxiliary image as input and synthesise a single reconstructed image, thus lacking the ability to handle long temporal auxiliary images, i.e., Dual-In Single-Out. In contrast, our proposed method leverages multi-temporal images with missing data as input and outputs reconstructed multi-temporal images in parallel, i.e., Multi-In Multi-Out. The simulated experiments utilized the mean peak signal-to-noise ratio (mPSNR) and the mean structural similarity index (mSSIM) for all spectral bands [6] as evaluation metrics. Furthermore, we repeated each experiment 5 times and used the average results as the final evaluation metric.
IV-A3 Implement Detail
For deep learning-based models, we utilized the Adam optimizer with parameters ($\beta_{1}=0.9$, $\beta_{2}=0.999$) for training, and set the batch size to 8. The initial learning rate was set to $4\times 10^{-4}$, with a decay schedule of every 100 epochs, and an early stopping strategy of 30 epochs on the validation set to prevent overfitting. The proposed MS2TAN was implemented with PyTorch 1.12 framework and trained with a NVIDIA GeForce RTX3090 24GB GPU on a Ubuntu 20.04 environment. Additionally, the test code of MS2TAN is openly available on the Github at https://github.com/zzaiyan/MS2TAN.
IV-B Simulated Experiment
Figure 7: Experimental results of Simulated ETM+ SLC-off Restoration based on Research Area S1. (a) Ground truth (March 19, 2002). (b) Simulated ETM+ SLC-off. (c) LLHM+LPRM. (d) WLR+LPRM. (e) STS-CNN. (f) Ours.
Figure 8: Experimental results of Simulated ETM+ SLC-off Restoration based on Research Area S2. (a) Ground truth (April 15, 2003). (b) Simulated ETM+ SLC-off. (c) LLHM+LPRM. (d) WLR+LPRM. (e) STS-CNN. (f) Ours.
IV-B1 Simulated ETM+ SLC-off Restoration Experiment
For the land satellite ETM+ SLC-Off problem, we conducted simulated experiments with synthesized missing data in research areas S1 and S2, as illustrated in Fig. 7-8. In these figures, (a) and (b) represent the original image and the simulated SLC-off image, respectively, while (c)-(f) show the simulated recovery results using four methods—LLHM, WLR, STS-CNN, and the proposed MS2TAN. To provide a clearer view of the recovery results, we present magnified details next to each reconstruction. As shown in Fig. 7-8 (c)-(e), all compared methods generate discontinuous fine features to some extent. This is because the temporal auxiliary images cannot fully cover the missing regions, thus requiring LLHM and WLR algorithms to rely on LPRM to fill the remaining gaps, resulting in highly blurred noise bands in the reconstructed results. Although the STS-CNN model based on an end-to-end strategy can repair these gaps, it still cannot accurately restore the original features. In contrast, the proposed MS2TAN, utilizing a Masked Spatial-Temporal Attention mechanism and integrating a “Pixel-Structure-Perception” Multi-Objective Optimization approach, achieves better restoration of texture and structural details, as shown in Fig. 7-8 (f). Moreover, this method outperforms state-of-the-art methods in the quantitative evaluations presented in Table II, validating its strong capability in spatio-temporal integrated data utilization.
Figure 9: Experimental results of Simulated Thick Cloud Removal based on Research Area S3. (a) Ground truth (October 17, 2003). (b) Simulated Thick Cloud. (c) LLHM. (d) WLR. (e) STS-CNN. (f) Ours.
Figure 10: Experimental results of Simulated Thick Cloud Removal based on Research Area S4. (a) Ground truth (January 22, 2014). (b) Simulated Thick Cloud. (c) LLHM. (d) WLR. (e) STS-CNN. (f) Ours.
TABLE II: Quantitative Evaluation Result of the Simulated SLC-off Restoration in Research Area S1 and S2
  Dataset Area S1 Area S2
Metric mPSNR($\uparrow$) mSSIM($\uparrow$) mPSNR($\uparrow$) mSSIM($\uparrow$)
  LLHM [21] 32.424 0.8947 32.908 0.9049
WLR [24] 34.374 0.9179 34.498 0.9214
STS-CNN [37] 37.118 0.9446 37.761 0.9479
$\textbf{MS}^{2}\textbf{TAN}$ (Ours) 38.255 0.9548 38.963 0.9555
IV-B2 Simulated Thick Cloud Removal Experiment
Similar to the simulated experiments of ETM+ SLC-off, we simulated thick cloud removal tasks in research areas S3 and S4, and reconstructed the data using the same four methods, resulting in the outcomes shown in Fig. 9-10. Among them, LLHM exhibits severe texture discontinuities, while WLR produces numerous noise points. This is because these simple models struggle to handle highly complex nonlinear relationships between different times. In comparison, the STS-CNN model can more accurately reconstruct the texture and structural features of missing parts, but exhibits noticeable spectral transition phenomena at the edges of missing value regions. On the other hand, the proposed MS2TAN method introduces missing value hint tensors into the learning process, enhancing the model’s capability to handle irregularly distributed missing values. Ultimately, it performs well in maintaining consistency in texture, structure, and spectral properties, achieving the best metrics in the quantitative evaluation presented in Table III.
TABLE III: Quantitative Evaluation Result of the Simulated Thick Cloud Removal in Research Area S3 and S4
  Dataset Area S3 Area S4
Metric mPSNR($\uparrow$) mSSIM($\uparrow$) mPSNR($\uparrow$) mSSIM($\uparrow$)
  LLHM [21] 34.775 0.9439 34.714 0.9370
WLR [24] 39.003 0.9639 38.261 0.9614
STS-CNN [37] 39.515 0.9711 38.753 0.9662
$\textbf{MS}^{2}\textbf{TAN}$ (Ours) 39.802 0.9753 39.272 0.9693
Figure 11: Experimental results of Real ETM+ SLC-off Restoration based on Research Area S5. (a) SLC-off Image 1 (October 23, 2011). (b) SLC-off Image 2 (November 8, 2011). (c) LLHM+LPRM. (d) WLR+LPRM. (e) STS-CNN. (f) Ours.
Figure 12: Experimental results of Real Thick Cloud Removal based on Research Area S6. (a) Image with clouds (August 5, 2007). (b) Temporal image without clouds (July 20, 2007). (c) LLHM. (d) WLR. (e) STS-CNN. (f) Ours.
IV-C Real Data Experiment
IV-C1 SLC-off Restoration Experiment
The experimental results of real SLC-off images in research area S5 are shown in Fig. 11. In this figure, (a) and (b) depict two ETM+ SLC-off images observed on October 23, 2011, and November 8, 2011, respectively, while (c)-(h) display the outputs of four algorithms. Similar to the simulated SLC-off reconstruction experiment, single temporal auxiliary images cannot fully cover the gaps. Thus, after LLHM and WLR processing, the remaining gaps are filled using the LPRM algorithm. STS-CNN can directly repair the entire image but exhibits discontinuous texture features similar to the simulated experiment. Thanks to the “Pixel-Structure-Perception” Multi-Objective Joint Optimization method, the proposed MS2TAN method preserves more structural details and produces more consistent reconstruction results.
IV-C2 Thick Cloud Removal Experiment
The experimental results of thick cloud coverage reconstruction in region S6 are shown in Fig. 12. In Fig. 12 (a), an image containing thick clouds is presented, while (b) depicts the temporal auxiliary image. Four algorithms identical to those used in the SLC-off reconstruction experiment were employed to reconstruct the images, yielding results displayed in (c)-(f). It can be observed in Fig. 12 (a)-(b) that for extensive cloud coverage and cloud shadows, the results of LLHM and WLR models suffer from significant loss of texture and structural features, presenting as blurry patches. This is attributed to the inconsistency between the reconstructed texture details in the cloud region and the surrounding non-cloud areas, which the aforementioned models fail to fit due to their incapability to capture the nonlinear relationship between different temporal data. In comparison, the STS-CNN method exhibits better performance in texture restoration, yet still shows noticeable spectral discontinuities at the boundaries of missing regions. On the other hand, the proposed MS2TAN method fully preserves texture and structural details with less spectral distortion compared to STS-CNN, resulting in high-quality thick cloud removal images.
IV-D Validation Study
IV-D1 Masked Spatial-Temporal Attention
To validate the effectiveness of the proposed Masked Spatial-Temporal Attention (MSTA) mechanism, we compared models with and without the adoption of this mechanism, and plotted their mSSIM iteration curves on the validation set, as shown in Fig. 13. It can be observed that with the assistance of the attention mask mechanism, the model exhibits faster convergence speed and higher fitting accuracy. This confirms that the attention mask mechanism effectively enhances the model’s feature extraction capability and training speed.
Figure 13: Ablation experiments for the proposed MSTA mechanism. The blue and red curves respectively depict the variations of mSSIM with training rounds for models with and without the adoption of the attention mask mechanism in Area S1.
IV-D2 Multi-scale Restoration
The MS2TAN employs a multi-scale restoration strategy to process input images. To validate its effectiveness, we take the SLC-off reconstruction task as an example and compare the intermediate outputs of each layer. The results are shown in Fig. 14. In the figure, the missing regions of the input image are filled with zeros. After Coarse Embedding, MFE, and Unembedding, a coarse repair result is obtained. This result reconstructs most of the texture and structural features and balances the significant distribution differences between the missing regions and other areas in the input. However, there are still spectral discontinuities at the edges of the missing regions. After several layers of restoration, the final refined repair result is obtained. This result refines the previous intermediate results, with clearer textures, consistent spectra, and smoother integration of the repair area boundary with other areas.
Figure 14: Ablation experiments for multi-scale restoration. (a)-(c) represent observations from three different times, processed in parallel by MS2TAN after composing the time-series remote sensing images. From top to bottom, they represent: input with missing values, coarse result, and fine result.
IV-D3 Gains from long time-series image inputs
A highlight of the proposed model is its capability to utilize longer time-series images for assistance in restoration, leading to more accurate reconstruction results. We designed experiments to compare different lengths of time-series inputs to quantitatively evaluate the accuracy gains from longer time-series inputs. The experimental results are shown in Fig. 15. From Fig. 15, it can be observed that as the length of the input time series increases from 2, there is a significant improvement in mPSNR and mSSIM metrics. However, when the time series length reaches 5 or longer, the change in accuracy becomes less significant. Therefore, longer time-series inputs can significantly improve the accuracy of missing value reconstruction, but there is a certain limit. Hence, we can find a balance point between reconstruction accuracy and algorithm flexibility.
Figure 15: Analysis of the input time-series length. The left and right tables record the variations of mPSNR and mSSIM with the length of the time-series, respectively, and distinguish the results of four research areas with different colors.
V Conclusion
In this paper, we proposed a novel method for time-series remote sensing image reconstruction based on a Multi-scale Masked Spatial-Temporal Attention Network. The proposed method reconstructs various missing value scenarios, including the Landsat ETM+ SLC-off problem and removal of thick cloud coverage in remote sensing images. Compared to existing methods, the proposed method improves the efficiency of spatio-temporal information utilization, enabling the mutual assistance of multiple degraded time-series images for reconstruction and achieving higher restoration accuracy through more auxiliary information. Additionally, the application of the MSTA mechanism and Multi-Objective Joint Optimization further enhances the texture and structural consistency of the reconstruction results. Experimental results demonstrate significant improvements in visual effects and restoration accuracy compared to mainstream methods.
Although the proposed method performs well in reconstructing the ETM+ SLC-off problem and thick cloud removal, there are still some inevitable limitations. For example, the model needs to simulate various shapes of cloud-like missing regions in the dataset and learn the characteristics of missing value distribution extraction sufficiently to achieve better cloud removal effects, which requires more training time for the model. In our future research, we can incorporate network components pre-trained on large-scale datasets into the model to improve its convergence speed.
Acknowledgments
This work was supported in part by the National Key R&D Program of China (Project No. 2022YFC3800700); in part by the National Natural Science Foundation of China (No. U21A2013).

### v2

@@
 Abstract
-Due to factors such as thick cloud cover and sensor limitations, remote sensing images often suffer from significant missing data, resulting in incomplete time-series information.
-Existing methods for imputing missing values in remote sensing images do not fully exploit spatio-temporal auxiliary information, leading to limited accuracy in restoration.
-Therefore, this paper proposes a novel deep learning-based approach called MS2TAN (Multi-scale Masked Spatial-Temporal Attention Network), for reconstructing time-series remote sensing images.
-Firstly, we introduce an efficient spatio-temporal feature extractor based on Masked Spatial-Temporal Attention (MSTA), to obtain high-quality representations of the spatio-temporal neighborhood features in the missing regions.
-Secondly, a Multi-scale Restoration Network consisting of the MSTA-based Feature Extractors, is employed to progressively refine the missing values by exploring spatio-temporal neighborhood features at different scales.
-Thirdly, we propose a “Pixel-Structure-Perception” Multi-Objective Joint Optimization method to enhance the visual effects of the reconstruction results from multiple perspectives and preserve more texture structures.
-Furthermore, the proposed method reconstructs missing values in all input temporal phases in parallel (i.e., Multi-In Multi-Out), achieving higher processing efficiency.
-Finally, experimental evaluations on two typical missing data restoration tasks across multiple research areas demonstrate that the proposed method outperforms state-of-the-art methods with an improvement of 0.40dB/1.17dB in mean peak signal-to-noise ratio (mPSNR) and 3.77‰/9.41‰ in mean structural similarity (mSSIM), while exhibiting stronger texture and structural consistency.
+Remote sensing images often suffer from substantial data loss due to factors such as thick cloud cover and sensor limitations.
+Existing methods for imputing missing values in remote sensing images fail to fully exploit spatiotemporal auxiliary information, which restricts the accuracy of their reconstructions.
+To address this issue, this paper proposes a novel deep learning-based approach called MS2TAN (Multi-Scale Masked Spatial-Temporal Attention Network) for reconstructing time-series remote sensing images.
+First, we introduce an efficient spatiotemporal feature extractor based on Masked Spatial-Temporal Attention (MSTA) to capture high-quality representations of spatiotemporal neighborhood features surrounding missing regions while significantly reducing the computational complexity of the attention mechanism.
+Second, a Multi-Scale Restoration Network composed of MSTA-based Feature Extractors is designed to progressively refine missing values by exploring spatiotemporal neighborhood features at different scales.
+Third, we propose a “Pixel-Structure-Perception” Multi-Objective Joint Optimization method to enhance the visual quality of the reconstructed results from multiple perspectives and to preserve more texture structures.
+Finally, quantitative experimental results under multi-temporal inputs on two public datasets demonstrate that the proposed method outperforms competitive approaches, achieving a 9.76%/9.30% reduction in Mean Absolute Error (MAE) and a 0.56 dB/0.62 dB increase in Peak Signal-to-Noise Ratio (PSNR), along with stronger texture and structural consistency.
+Ablation experiments further validate the contribution of the core innovations to imputation accuracy.
 Index Terms:
@@
 Researchers have proposed various methods to recover missing data in remote sensing images.
-Early approaches to missing value restoration can be broadly categorized into three types [1]: spatial-based, temporal-based, and spatio-temporal-based data recovery.
+Early approaches to missing value restoration can be broadly categorized into three types [1]: spatial-based, temporal-based, and spatiotemporal-based data recovery.
 These methods have shown promising results in specific scenarios with low resolution and low missing rates.
@@
 In recent years, the rapid advancement of deep learning theory and computer hardware [2] has led to significant progress in remote sensing image restoration using deep learning-based methods.
-These methods can be broadly categorized into two types: CNN-based [3] and Transformer-based [4] approaches.
+These methods can be broadly categorized into two types: CNN-based [3] and ViT-based [4] approaches.
 Compared to traditional statistical models, CNNs exhibit strong non-linear expressive power, allowing for efficient extraction of spatial features from remote sensing data and significantly improving the accuracy of image restoration.
 However, these structures struggle to fully exploit time-series information, resulting in a bottleneck in restoration accuracy.
-Transformers, based on self-attention mechanisms, possess a global receptive field, enabling comprehensive and efficient utilization of both temporal sequences and images [5] to enhance reconstruction results [6].
+ViTs, based on self-attention mechanism [5], possess a global receptive field, enabling comprehensive and efficient utilization of both temporal sequences and images to enhance reconstruction results [6].
 However, due to the high resolution and long time-series of remote sensing images, token sequences become excessively long, leading to extremely high complexity in attention computations.
-To efficiently mine spatio-temporal information in remote sensing images, we apply self-attention mechanism separately in the temporal and spatial dimensions and alternate between the two [7], greatly reducing computational complexity.
-To address the problem of significant distributional differences between missing and non-missing values in remote sensing images, we apply missing values mask and diagonal mask to the attention matrix [8], proposing Masked Spatial-Temporal Attention (MSTA) to enhance the expressive power of spatio-temporal attention and optimize the spectral discrepancy at the transition regions.
-For full exploitation of the spatio-temporal neighborhood features at different scales [9], we further propose a Multi-scale Restoration Network.
+To efficiently mine spatiotemporal information in remote sensing images, we apply self-attention mechanism separately in the temporal and spatial dimensions and alternate between the two [7], greatly reducing computational complexity.
+To address the problem of significant distributional differences between missing and non-missing values in remote sensing images, we apply missing values mask and diagonal mask to the attention matrix [8], proposing Masked Spatial-Temporal Attention (MSTA) to enhance the expressive power of spatiotemporal attention and optimize the spectral discrepancy at the transition regions.
+For full exploitation of the spatiotemporal neighborhood features at different scales [9], we further propose a Multi-scale Restoration Network.
 The network consists of MSTA-based Feature Extractors with different embedding scales, which progressively refine the reconstruction of missing information from coarse to fine granularity levels, achieving higher restoration accuracy [10].
 For model training, we propose a “Pixel-Structure-Perception” Multi-Objective Joint Optimization method, using pixel-wise loss as the basic loss for the restoration task and incorporating structural loss [11] and perceptual loss [12] to optimize the model’s results from the perspectives of structure, texture, shape, and spatial relations, thereby achieving high-quality image inpainting.
-Finally, we conducted simulated and real data experiments to validate the performance of the proposed method for the ETM+ SLC-off and thick cloud removal tasks.
-The results from six research areas indicate that our method outperforms competitors.
-In addition, a validation study demonstrates the effectiveness of the core components.
+Finally, we performed both quantitative and qualitative experiments on two datasets, comparing our approach with several mainstream methods.
+Additionally, we carried out ablation studies on the key innovations and analyzed the trade-offs between effectiveness and efficiency across models of different sizes.
 In summary, our main contributions are as follows:
   1. 1.
-    We propose a deep learning-based approach called MS2TAN, for repairing missing information in time-series remote sensing images.
-Our method utilizes a Multi-scale Restoration Network with MSTA-based Feature Extractors to learn the non-linear mapping between missing and complete data.
-Both simulated and real data experiments show that proposed MS2TAN outperforms state-of-the-art methods across multiple research areas.
+    We propose a deep learning-based method MS2TAN for reconstructing missing data in time series remote sensing images.
+Our method utilizes a multi-scale restoration network to learn an end-to-end mapping between incomplete and complete image sequences.
+MS2TAN achieved higher restoration accuracy than mainstream methods in quantitative experiments, and showed better visual effects in eliminating gaps in real data.
   2. 2.
-    To address the challenges of high resolution and long temporal sequences in remote sensing data, we introduce the Masked Spatial-Temporal Attention (MSTA) mechanism.
-The MSTA efficiently extracts spatio-temporal features, improves the utilization of spatio-temporal contextual information, reduces color transition artifacts at the boundaries of missing value regions, and significantly reduces the time complexity of self-attention.
+    To address the challenges of high resolution and long time series in remote sensing data, we introduce the Masked Spatial-Temporal Attention (MSTA) mechanism.
+MSTA effectively extracts spatiotemporal features, improves the utilization of spatiotemporal context information, reduces color transition artifacts at the boundaries of missing value areas, and significantly reduces the computational complexity of self-attention.
   3. 3.
@@
 The network architecture and methodology details of our proposed model are presented in Section III.
-In Section IV, we showcase the results of missing data reconstruction in both simulated and real data experiments, compare them with mainstream methods, and validate the effectiveness of each component through validation study.
+In Section IV, we showcase the results of missing data reconstruction in both quantitative and qualitative experiments, compare them with mainstream methods, and validate the effectiveness of each component through validation studies.
 Finally, our conclusions and future prospects are summarized in Section V.
@@
 II-A Traditional methods
-Early research on the restoration of missing data in remote sensing images can be roughly divided into three categories: spatial-based, temporal-based, and spatio-temporal hybrid methods.
+Early research on the restoration of missing data in remote sensing images can be roughly divided into three categories: spatial-based, temporal-based, and spatiotemporal hybrid methods.
 II-A1 Spatial-based Methods
@@
 However, the differences between different time phases restrict the application of these methods.
-II-A3 Spatio-temporal-based Methods
-To overcome these limitations, spatio-temporal-based methods integrate the spatial and temporal correlations to reconstruct missing data under various conditions.
-For instance, improved nearest neighbor pixel interpolation methods [27], methods based on spatio-temporal Markov random field model [28], spatio-temporal weighted regression model [29], methods based on group sparse representation [30], and methods based on low-rank tensor decomposition [31, 32] have been proposed.
+II-A3 Spatiotemporal Hybrid Methods
+To overcome these limitations, spatiotemporal hybrid methods integrate the spatial and temporal correlations to reconstruct missing data under various conditions.
+For instance, improved nearest neighbor pixel interpolation methods [27], methods based on spatiotemporal Markov random field model [28], spatiotemporal weighted regression model [29], methods based on group sparse representation [30], and methods based on low-rank tensor decomposition [31, 32] have been proposed.
 However, most of these methods rely on linear models and struggle to handle complex and intricate scenes.
@@
 Zhang et al.
-[37] proposed a CNN-based Spatio-Temporal Spectra (STS-CNN) framework, which was further developed into a progressive spatio-temporal patch grouping framework [38].
+[37] proposed a CNN-based spatiotemporal Spectra (STS-CNN) framework, which was further developed into a progressive spatiotemporal patch grouping framework [38].
 CNNs that incorporate temporal inputs through channel-wise concatenation [39] merge feature maps of target images and temporal images, introducing auxiliary information for missing data restoration.
-However, CNNs lack a true understanding of time series, which hinders the efficient integration of temporal and spatial information in these methods and limits the utilization of long temporal sequences as auxiliary information, ultimately affecting the restoration accuracy.
-II-B2 Transformer-based methods
-Transformer [4] employs self-attention mechanism to process text sequences and capture global contextual information.
-It has been proven that Transformer-based methods, such as Visual Transformer (ViT) [5], also exhibit powerful performance in the field of image processing.
-The sequence modeling capability of self-attention mechanism can be leveraged to assist in high-precision restoration of missing values in remote sensing images.
+Stucker et al.
+[40] used temporal self-attention for CNN’s feature map sequences in conjunction with U-Net to repair temporal images.
+However, CNNs lack a true understanding of time-series, which hinders the efficient integration of temporal and spatial information in these methods and limits the utilization of long temporal sequences as auxiliary information, ultimately affecting the restoration accuracy.
+II-B2 ViT-based methods
+Visual Transformer (ViT) [4] has shown excellent performance in many tasks in the vision domain.
 Xu et al.
@@
 [43] used spatial attention and channel attention to remove cloud cover in the images.
-Li et al.
-[43] introduced SAR auxiliary data and synthesized cloud-free images using a Transformer-based GAN.
-However, the spatial attention used in ViT requires computing the correlations between all pairs of patches.
+However, the self-attention used in ViT requires computing the correlations between all pairs of patches (including all times).
 In the context of processing time-series remote sensing images, the images often have high resolutions and long time-series, resulting in a large number of tokens and extremely high computational complexity.
@@
 We further improved the separated spatial-temporal attention by introducing missing value masks and diagonal masks [8], resulting in Masked Spatial-Temporal Attention (MSTA), which exhibits superior performance in the task of missing value restoration.
-Compared to CNN-based methods and original ViT methods, MSTA enables efficient processing of spatio-temporal information and leverages long-term temporal information to assist in missing value reconstruction, leading to more precise reconstruction results.
+Compared to CNN-based methods and original ViT methods, MSTA enables efficient processing of spatiotemporal information and leverages long-term temporal information to assist in missing value reconstruction, leading to more precise reconstruction results.
 Figure 1: The overall flowchart of the proposed method consists of two main components: a Multi-scale Restoration Network (named MS2TAN) with MSTA-based Feature Extractors, and a “Pixel-Structure-Perception” Multi-Objective Joint Optimization method.
@@
 III-A1 Problem Definition
-The original time-series remote sensing image with missing values can be represented as $X\in\mathbb{R}^{T\times C\times H\times W}$, where $T$ denotes the length of the time series, $C$ represents the number of channels, and $H$ and $W$ denote the spatial dimensions of the region.
+The original time-series remote sensing image with missing values can be represented as $X\in\mathbb{R}^{T\times C\times H\times W}$, where $T$ denotes the length of the time-series, $C$ represents the number of channels, and $H$ and $W$ denote the spatial dimensions of the region.
 $X_{(t,c,i,j)}$ denotes the value of channel $c$ at position $(i,j)$ at time $t$.
-To account for the missing values, we introduce the missing value hint tensor $M\in\mathbb{R}^{T\times C\times H\times W}$, which is defined as shown in Equation (1).
-In practical applications, the missing value hint tensor $M$ can be obtained from QA band or cloud detection algorithms such as Fmask [44].
+To account for the missing values, we introduce the missing value hint tensor $M\in\mathbb{R}^{T\times C\times H\times W}$, which is defined as shown in Eq.
+(1).
+In practical applications, the missing value hint tensor $M$ can be obtained from QA band or cloud detection algorithms such as Fmask [44] and S2Cloudless [45].
 $M_{(t,c,i,j)}=\left\{\begin{matrix}0&\text{if}~{}X_{(t,c,i,j)}~{}\text{is~{}missing},\\ 1&\text{if}~{}X_{(t,c,i,j)}~{}\text{is~{}observed}.\end{matrix}\right.$ (1)
@@
 Here, the output $\widetilde{Y}$ of the model is the initial reconstruction result, and $\odot$ denotes the Hadamard product.
-By replacing the observed real values in $X$ with $\widetilde{Y}$, we obtain the final reconstruction result $\widetilde{Y}_{\mathrm{out}}$ as shown in equation (3).
+By replacing the observed real values in $X$ with $\widetilde{Y}$, we obtain the final reconstruction result $\widetilde{Y}_{\mathrm{out}}$ as shown in Eq.
+(3).
 $\widetilde{Y}=\operatorname{Model}(X,M)$ (2)
@@
 Here, $T$ represents the time-series length, $N$ represents the number of patches in a single remote sensing image, and $d_{\mathrm{emb}}$ represents the dimension of the token vector corresponding to each patch.
-Subsequently, the position encoding introduces spatial-temporal positional semantics, followed by $L$ layers of MSTA units, and ultimately produces a sequence with incorporated spatio-temporal correlation features.
+Subsequently, the position encoding introduces spatial-temporal positional semantics, followed by $L$ layers of MSTA units, and ultimately produces a sequence with incorporated spatiotemporal correlation features.
 III-B1 Masked Self-Attention
@@
 The attention score between Q and K is computed using Scaled Dot-Product, followed by the application of the Softmax function to obtain attention weights A.
-The final output is the attention-weighted V, as shown in Equation 4.
+The final output is the attention-weighted V, as shown in Eq.
+4.
 $\begin{gathered}\mathrm{H}=\mathrm{A}\mathrm{V}=\mathrm{Softmax}\left(\frac{\mathrm{Q}\cdot\mathrm{K}^{\top}}{\sqrt{d_{k}}}\right)\cdot\mathrm{V}\\ \mathrm{where}\ \mathrm{[Q,K,V]}=X\cdot W_{\mathrm{QKV}}.\end{gathered}$ (4)
@@
   • The diagonal mask sets the diagonal of the attention matrix as $\infty$- ∞, preventing each step from attending to itself and forcing it to rely on the other $TN-1$ steps for inference.
-This helps capture the spatio-temporal feature correlation in high-dimensional space.
-More specifically, for the attention score matrix $\mathrm{Sc}\in\mathbb{R}^{TN\times TN}$, the attention mask operation $\mathrm{ApplyMask}$ is defined as shown in Equation (5).
+This helps capture the spatiotemporal feature correlation in high-dimensional space.
+More specifically, for the attention score matrix $\mathrm{Sc}\in\mathbb{R}^{TN\times TN}$, the attention mask operation $\mathrm{ApplyMask}$ is defined as shown in Eq. (5).
 $\begin{gathered}\left.\mathrm{ApplyMask}\left(\mathrm{Sc}\right)_{\left(i,j\right)}=\left\{\begin{array}[]{ll}-\infty&\mathrm{Mask}(i,j)=1,\\ \mathrm{Sc}_{\left(i,j\right)}&\mathrm{Mask}(i,j)\neq 1.\end{array}\right.\right.\\ \mathrm{Mask}(i,j)=1\ \textbf{only if}\ (\mathrm{MR}(E(i))>C_{\mathrm{max}}\ \textbf{or}\ i=j)\end{gathered}$ (5)
 Here, $Mask()$ checks whether the mask condition is satisfied, $\mathrm{MR}(e)$ represents the missing rate of the corresponding patch $e$ (which can be calculated from the hint tensor $M$), and $C_{\mathrm{max}}$ is a hyperparameter that controls the maximum allowable missing rate.
-Applying Equation (5) to the attention scores in Equation (4), we obtain the expression for a single masked attention head, as shown in Equation (6).
-Finally, by linearly projecting and combining multiple attention heads with the residual connection, we obtain the output of the masked self-attention, as shown in Equation (7).
-$\mathrm{H}^{(i)}=\mathrm{A}^{\prime}\mathrm{V}=\mathrm{Softmax}\left(\mathrm{ApplyMask}\left(\frac{\mathrm{Q}\cdot\mathrm{K}^{\top}}{\sqrt{d_{k}}}\right)\right)\cdot\mathrm{V}$ (6)
+Applying Eq.
+(5) to the attention scores in Eq.
+(4), we obtain the expression for a single masked attention head, as shown in Eq.
+(6).
+Finally, by linearly projecting and combining multiple attention heads with the residual connection, we obtain the output of the masked self-attention, as shown in Eq.
+(7).
+$\mathrm{H}^{(i)}=\mathrm{A}^{\prime}\mathrm{V}=\mathrm{Softmax}\left(\mathrm{ApplyMask}\left(\frac{\mathrm{Q}\cdot\mathrm{K}^{\top}}{\sqrt{d_{\text{qkv}}}}\right)\right)\cdot\mathrm{V}$ (6)
 $\mathrm{MaskedAttn}(X,M)=\mathrm{Proj}\left(\left[\mathrm{H}^{(1)},\mathrm{H}^{(2)},\cdots,\mathrm{H}^{(h)}\right]\right)+X$ (7)
-In Equation (7), $\mathrm{A}^{\prime}$ denotes the attention weights post mask application, $\mathrm{MaskedAttn}()$ represents the masked self-attention operation, $\mathrm{Proj}$ stands for projection head, and $h$ denotes the number of attention heads.
-Figure 3: Schematic of MSTA.
-Taking the blue patch as an example, patches at the same spatial position are considered as temporal auxiliaries, while patches at the same time are regarded as spatial auxiliaries.
-Additionally, patches with excessively high missing rates will be excluded from the receptive field to avoid the influence of distributional differences on model learning.
+In Eq.
+(7), $\mathrm{A}^{\prime}$ denotes the attention weights post mask application, $\mathrm{MaskedAttn}()$ represents the masked self-attention operation, $\mathrm{Proj}()$ stands for projection head, and $h$ denotes the number of attention heads.
 III-B2 Masked Spatial-Temporal Attention
-Masked Spatial-Temporal Attention (MSTA) captures the spatio-temporal correlations within the input sequence, with its masking mechanism and receptive field depicted in Fig.
-3.
-Specifically, MSTA comprises Masked Temporal Attention (MTA), Masked Spatial Attention (MSA), Layer Normalization (LN), and Feed-Forward Network (FFN).
+Masked Spatial-Temporal Attention (MSTA) is used to capture the spatiotemporal correlations within the input sequence, and consists of Masked Temporal Attention (MTA), Masked Spatial Attention (MSA), Layer Normalization (LN), and Feed-Forward Network (FFN).
 These components are elaborated upon below.
 For an input sequence $E\in\mathbb{R}^{TN\times d_{\mathrm{emb}}}$, MTA first reshapes $E$ to place the time dimension and the feature dimension at the end, obtaining a sequence $e_{t}$ of length $T$ and $d_{\mathrm{emb}}$ dimensions.
-Subsequently, for each position in space, MaskedAttn operation is performed along the temporal direction to obtain the output of MTA, as shown in Equation (8).
+Subsequently, for each position in space, MaskedAttn operation is performed along the temporal direction to obtain the output of MTA, as shown in Eq.
+(8).
 $\begin{gathered}e_{t}=\mathrm{Reshape}(E)\in\mathbb{R}^{N\times T\times d_{\mathrm{emb}}}\\ \mathrm{MTA}(E,M)=\mathrm{Reshape}\left(\mathrm{MaskedAttn}\left(e_{t},M\right)\right)\end{gathered}$ (8)
-Similarly, after the corresponding dimension transformation of the input sequence, MaskedAttn operation is conducted along the spatial direction for each temporal image, yielding the output of MSA, as shown in Equation (9).
+Similarly, after the corresponding dimension transformation of the input sequence, MaskedAttn operation is conducted along the spatial direction for each temporal image, yielding the output of MSA, as shown in Eq.
+(9).
 $\begin{gathered}e_{s}=\mathrm{Reshape}(E)\in\mathbb{R}^{T\times N\times d_{\mathrm{emb}}}\\ \mathrm{MSA}(E,M)=\mathrm{Reshape}\left(\mathrm{MaskedAttn}\left(e_{s},M\right)\right)\end{gathered}$ (9)
-FFN consists of two linear layers separated by a ReLU activation function, with a residual connection established between the input and output. The expression for FFN is presented in Equation (10).
+FFN consists of two linear layers separated by a ReLU activation function, with a residual connection established between the input and output. The expression for FFN is presented in Eq. (10).
 $\mathrm{FFN}\left(X\right)=\mathrm{Linear}\left(\mathrm{ReLU}\left(\mathrm{Linear}\left(X\right)\right)\right)+X$ (10)
-MSTA sequentially applies MTA and MSA to the input token sequence, exploring the spatio-temporal correlations among patches.
-Subsequently, FFN is used to fuse spatio-temporal features and introduce non-linear transformations, as shown in Equation (11).
+MSTA sequentially applies MTA and MSA to the input token sequence, exploring the spatiotemporal correlations among patches.
+Subsequently, FFN is used to fuse spatiotemporal features and introduce non-linear transformations, as shown in Eq.
+(11).
 $\begin{gathered}U=\operatorname{MTA}(\operatorname{LN}(E),M)\\ V=\operatorname{MSA}(\operatorname{LN}(U),M)\\ \operatorname{MSTA}(E,M)=\operatorname{FFN}\left(\operatorname{LN}\left(V\right)\right)\end{gathered}$ (11)
-Figure 4: The network structure diagram of proposed MS2TAN (Multi-Scale Masked Spatial-Temporal Attention Network), which includes components such as Embedding, MFE, Unembedding, and Observed Value Replacement.
+Figure 3: The network structure diagram of proposed Multi-scale Restoration Network (MS2TAN), which contains several residual-connected restoration modules consisting of Embedding, MFE, and Unembedding.
 III-B3 Positional Encoding
-In the Transformer, positional encoding is added to the input sequence to introduce spatio-temporal positional semantics, as shown in Equation (12), where $\operatorname{PosEnc}(pos,dim)$ represents the value of the $dim$-th dimension of the positional encoding for the $pos$-th element in the sequence.
+In the Transformer, positional encoding is added to the input sequences to introduce positional semantics, as depicted in Eq.
+(12).
+Here, $pos=t\times N+n\in\left[0,~{}TN-1\right]$ represents the spatiotemporal joint coordinate derived from the temporal index $t\in\left[0,~{}T-1\right]$ and spatial index $n\in\left[0,~{}N-1\right]$, where $N$ denotes the total number of patches in a single image.
+The function $\operatorname{PosEnc}(pos,dim)$ indicates the value of the $dim$-th dimension of the positional encoding for the $pos$-th patch in the image sequence.
 $\begin{gathered}\operatorname{PosEnc}(pos,2i)=\sin\left({pos}\cdot{10000^{-{2i}/{d_{\mathrm{emb}}}}}\right)\\ \quad\operatorname{PosEnc}(pos,2i+1)=\cos\left({pos}\cdot{10000^{-{2i}/{d_{\mathrm{emb}}}}}\right)\end{gathered}$ (12)
-After positional encoding, the input sequence will undergo $L$ layers of MSTA to explore deeper spatio-temporal features.
-Finally, the overall expression of the MFE module is shown in Equation (13), where the symbol L denotes stacking $L$ layers, and $\mathrm{MFE}()$ represents the MSTA-based Feature Extractor.
+After positional encoding, the input sequence will undergo $L$ layers of MSTA to explore deeper spatiotemporal features.
+Finally, the overall expression of the MFE module is shown in Eq.
+(13), where the symbol L denotes stacking $L$ layers, and $\mathrm{MFE}()$ represents the MSTA-based Feature Extractor.
 $\begin{gathered}\mathrm{MFE}\left(E,M\right)=\{\mathrm{MSTA}\left(E+\mathrm{PosEnc},M\right)\}^{L}\end{gathered}$ (13)
@@
 The MS2TAN (Multi-Scale Masked Spatial-Temporal Attention Network) is a multi-scale restoration network composed of $S$ residual-connected restoration modules and observed value replacement, as illustrated in Fig.
-4.
+3.
 Each module consists of Patch Embedding, MFE, and Patch Unembedding.
-These restoration modules can capture the spatio-temporal correlations of pixels at different scales to predict missing values and ultimately replace observed values as outputs.
+These restoration modules can capture the spatiotemporal correlations of pixels at different scales to predict missing values and ultimately replace observed values as outputs.
 Figure 4: Illustration of Embedding and Unembedding processes. Embedding represents a series of transformations from left to right, and Unembedding represents the reverse.
 III-C1 Embedding and Unembedding
-The input to MS2TAN consists of two components: the image sequence $X_{\mathrm{in}}$ and the missing value hint tensor $M$.
-To incorporate the distributional properties of the missing values into the model learning process, $X_{\mathrm{in}}$ and $M$ are concatenated along the channel dimension to yield the merged input $X_{c}=\mathrm{Concat}(X_{\mathrm{in}},M)$, where $X_{c}=\left[x_{c}^{1},x_{c}^{2},\cdots,x_{c}^{T}\right]\in\mathbb{R}^{T\times C^{\prime}\times H\times W}$, where $C^{\prime}=C+1$ is the total number of channels after concatenation.
-Following the approach of Vision Transformer (ViT), each image is partitioned into a series of patches, with the size of each patch determined by the hyperparameter $P$.
+Following the approach of ViT, each image is partitioned into a series of patches, with the size of each patch determined by the hyperparameter $P$.
 For each temporal image, $N=HW/P^{2}$ patches are obtained.
-These patches are then aggregated and reorganized to form a patch sequence of length $TN$, with each patch represented as a $C^{\prime}P^{2}$-dimensional vector.
-Equation (14) delineates the processing of the $t$-th temporal image $x_{c}^{t}$.
-$\begin{gathered}A^{t}=\left[p_{1}^{t},p_{2}^{t},\cdots,p_{N}^{t}\right]=\operatorname{Patchify}(x_{c}^{t},P)\\ \text{where}\ x_{c}^{t}\in\mathbb{R}^{C^{\prime}\times H\times W}\ \text{and}\ A^{t}\in\mathbb{R}^{N\times C^{\prime}P^{2}}\end{gathered}$ (14)
+These patches are then aggregated and reorganized to form a patch sequence of length $TN$, with each patch represented as a $CP^{2}$-dimensional vector.
+Equation (14) delineates the processing of the $t$-th temporal image $X_{\mathrm{in}}^{t}$.
+$\begin{gathered}A^{t}=\left[p_{1}^{t},p_{2}^{t},\cdots,p_{N}^{t}\right]=\operatorname{Patchify}(x_{\mathrm{in}}^{t},P)\\ \text{where}\ x_{\mathrm{in}}^{t}\in\mathbb{R}^{C\times H\times W}\ \text{and}\ A^{t}\in\mathbb{R}^{N\times CP^{2}}\end{gathered}$ (14)
 Next, a linear layer is utilized to project the vectors corresponding to each patch into a high-dimensional representation space.
-The images from different time steps are then aggregated to obtain the complete token sequence $E$, as depicted in Equation (15).
+The images from different time steps are then aggregated to obtain the complete token sequence $E$ :
 $\begin{gathered}E^{t}=\mathrm{Linear}(A^{t})\in\mathbb{R}^{N\times d_{\mathrm{emb}}}\\ E=\left[E^{1},E^{2},\cdots,E^{T}\right]\in\mathbb{R}^{TN\times d_{\mathrm{emb}}}\end{gathered}$ (15)
-The process of transforming $X_{\mathrm{in}}$ and $M$ step by step to obtain $E$ as described above is referred to as Embedding, which can be represented by Equation (16).
-Conversely, if this operation is reversed, i.e., $E$ is restored to a patch sequence and then rearranged into an image, it is referred to as Unembedding, as shown in Equation (17).
+The process of transforming $X_{\mathrm{in}}$ and $M$ step by step to obtain $E$ as described above is referred to as Embedding, which can be represented by Eq.
+(16).
+Conversely, if this operation is reversed, i.e., $E$ is restored to a patch sequence and then rearranged into an image, it is referred to as Unembedding, as shown in Eq.
+(17).
 Fig.
 4 illustrates the processes of Embedding and Unembedding on images.
-$E=\mathrm{Embedding}\left(\mathrm{Concat}\left(X_{\mathrm{in}},M\right),P\right)$ (16)
+$E=\mathrm{Embedding}\left(X_{\mathrm{in}},P\right)$ (16)
 $X_{\mathrm{out}}=\mathrm{Unembedding}\left(E,P\right)$ (17)
 III-C2 MS2TAN
-MS2TAN consists of components such as Embedding, MFE, and Unembedding, where the block size $P$ is an important hyperparameter, and the embedding scale of the $i$-th layer is denoted as $P^{\left(i\right)}$.
-The process of obtaining the embedding vector $\alpha$ from the input is described in Equation (18).
-Subsequently, MFE is utilized to mine deep spatio-temporal correlation features $\beta$, as shown in Equation (19).
-Finally, the features are decoded, unembedded back into image form, and added to the input, resulting in an intermediate result as depicted in Equation (20), where $\widetilde{Y}^{(i)}$ represents the $i$-th intermediate result.
-$\alpha=\mathrm{Embedding}\left(\mathrm{Concat}\left(\widetilde{Y}^{(i-1)},M\right),P^{\left(i\right)}\right)$ (18)
+MS2TAN consists of components such as Embedding, MFE, and Unembedding, where the patch size $P$ is an important hyperparameter, and the embedding scale of the $i$-th layer is denoted as $P^{\left(i\right)}$.
+The process of obtaining the embedding vector $\alpha$ from the input is described in Eq.
+(18).
+Subsequently, MFE is utilized to mine deep spatiotemporal correlation features $\beta$, as shown in Eq.
+(19).
+Finally, the features are decoded, unembedded back into image form, and added to the input, resulting in an intermediate result as depicted in Eq.
+(20), where $\widetilde{Y}^{(i)}$ represents the $i$-th intermediate result.
+$\alpha=\mathrm{Embedding}\left(\widetilde{Y}^{(i-1)},P^{\left(i\right)}\right)$ (18)
 $\beta=\mathrm{MFE}\left(\alpha,M\right)$ (19)
 $\widetilde{Y}^{(i)}=\mathrm{Unembedding}\left(\beta,P^{\left(i\right)}\right)+\widetilde{Y}^{(i-1)}$ (20)
-Let $\widetilde{Y}^{(0)}=X_{\mathrm{in}}$, and sequentially obtain $S$ intermediate results of the reconstruction.
-Eventually, the $S$-th result is taken as the output of MS2TAN, as shown in Equation (21).
+Let $\widetilde{Y}^{(0)}=X_{\mathrm{in}}$, and sequentially obtain $S$ intermediate results of the reconstruction. Eventually, the $S$-th result is taken as the output of MS2TAN, as shown in Eq. (21).
 […]
 III-C3 Observed Value Replacement
-For the original reconstruction results of this network, denoted as $\widetilde{Y}$, the observed values are replaced with their non-missing parts to obtain the final result $\widetilde{Y}_{\mathrm{out}}$, as shown in Equation (22).
-$\widetilde{Y}_{\mathrm{out}}=\widetilde{Y}\odot(1-M)+X\odot M$ (22)
+For the original reconstruction results of this network, denoted as $\widetilde{Y}$, the observed values are replaced with their non-missing parts to obtain the final result $\widetilde{Y}_{\mathrm{out}}$, as shown in Eq.
+(3).
+III-C4 Multi-scale Restoration Strategy
+Common ViT-based models use a fixed patch size, which has been shown to be inefficient in dense prediction tasks like image segmentation and restoration [46].
+In such tasks, smaller patch sizes often achieve better performance, but also come with greater computational costs.
+In order to balance the effect and efficiency, we apply a multi-scale restoration strategy, applying a larger patch size at a scale close to the input to obtain a coarse result, and applying a smaller patch size at a scale close to the output to obtain a fine result.
 III-D Multi-Objective Joint Optimization
@@
 Pixel-wise loss disregards the overall integrity of the image, treating the image as a collection of pixels, and comparing pixel by pixel to generate the image against the target image.
-It serves as the foundational loss for remote sensing image reconstruction tasks.
-Specifically, Equation (23) illustrates its calculation process.
-${\mathcal{L}}_{\textrm{pixel-wise}}=\frac{1}{TCHW}\left|\left|\eta-y\right|\right|_{2}^{2}$ (23)
+It serves as the foundational loss for image reconstruction tasks.
+Equation (22) illustrates its calculation process.
 The pixel-wise loss calculated here includes both missing and observed parts.
@@
 However, it does smooth the model output, making the model training process more stable, and hence is also included.
+${\mathcal{L}}_{\textrm{pixel-wise}}=\frac{1}{TCHW}\left|\left|\eta-y\right|\right|_{2}^{2}$ (22)
 III-D2 Structural Loss
-Pixel-wise reconstruction loss is commonly employed in various time-series and visual tasks but fails to consider the correlation between pixels.
-Structural reconstruction loss, based on the principle of structural similarity [45], measures the difference with the target image using the structural similarity index (SSIM), optimizing the visual consistency of the reconstruction results from the perspectives of structure, contrast, and luminance.
+Pixel-wise loss is commonly employed in various time-series and visual tasks but fails to consider the correlation between pixels.
+Structural loss uses structural similarity [47] to measure the difference from the target image to optimise the visual consistency of the reconstruction results in terms of structure, contrast and luminance, and has been shown to have better performance in image reconstruction tasks [48].
+Equation (23) specifies the computation process of structural loss, where $\mu_{\eta}$ and $\mu_{y}$ denote the means of $\eta$ and $y$ respectively, $\sigma_{\eta}$ and $\sigma_{y}$ denote the variances of $\eta$ and $y$ respectively, $\sigma_{\eta y}$ denotes the covariance between $\eta$ and $y$, and $C_{1}$ and $C_{2}$ are constants.
 ${\mathcal{L}}_{\mathrm{structural}}$ $=1-\mathrm{SSIM}\left(\eta,y\right)$ (23)
 $=1-\frac{\left(2{\mu_{\eta}}{\mu_{y}}+C_{1}\right)\left(2\sigma_{\eta y}+C_{2}\right)}{\left(\mu_{\eta}^{2}+\mu_{y}^{2}+C_{1}\right)\left(\sigma_{\eta}^{2}+\sigma_{y}^{2}+C_{2}\right)}$
-Equation (24) specifies the computation process of structural loss, where $\mu_{\eta}$ and $\mu_{y}$ denote the means of $\eta$ and $y$ respectively, $\sigma_{\eta}$ and $\sigma_{y}$ denote the variances of $\eta$ and $y$ respectively, $\sigma_{\eta y}$ denotes the covariance between $\eta$ and $y$, and $C_{1}$ and $C_{2}$ are constants.
 III-D3 Perceptual Loss
-Perceptual loss utilizes a pre-trained VGG16 network [46], as the feature extraction network to obtain perceptual feature vectors, followed by applying L2 loss to the feature vectors.
-The calculation of perceptual loss is shown in Equation (25), where $\psi\left(\right)$ represents the feature extraction network, and $d_{f}$ denotes the feature dimension.
-${\mathcal{L}}_{\mathrm{perceptual}}=\frac{1}{d_{f}}\left|\left|\psi\left(\eta\right)-\psi\left(y\right)\right|\right|_{2}^{2}$ (25)
+Perceptual loss [12] is widely used in tasks like super-resolution and image generation, where pre-trained CNNs are used as the feature network to extract perceptual features such as texture details, image style, etc.
+Specifically, we use a VGG-16 [49] with the fully connected layer removed as the feature network and the L2 loss of the feature map differences as the final loss.
+The calculation of perceptual loss is shown in Eq.
+(24), where $\psi\left(\right)$ denotes the feature network and $d_{\text{f}}$ denotes the feature dimension.
+${\mathcal{L}}_{\mathrm{perceptual}}=\frac{1}{d_{\text{f}}}\left|\left|\psi\left(\eta\right)-\psi\left(y\right)\right|\right|_{2}^{2}$ (24)
 III-D4 Multi-Objective Loss Function
-For each intermediate output $\widetilde{Y}^{(i)}$ of MS2TAN, the expression of the multi-objective loss function is shown in Equation (26), where $\lambda_{\mathrm{s}}$ and $\lambda_{\mathrm{p}}$ denote the weights of the structural loss and perceptual loss respectively.
-The final loss function considers the intermediate results of each layer, as shown in Equation (27).
-$\mathcal{L}^{\left(i\right)}\left(\eta,y\right)=\mathcal{L}_{\text{pixel-wise}}+{\lambda_{\mathrm{s}}\mathcal{L}}_{\mathrm{struct}}+{\lambda_{\mathrm{p}}\mathcal{L}}_{\mathrm{perceptual}}$ (26)
-$\mathcal{L}=\frac{1}{S}\sum_{i=1}^{m}\mathcal{L}^{\left(i\right)}\left(\widetilde{Y}^{(i)},Y\right)$ (27)
-Figure 6: The locations of the research areas used to evaluate the proposed methodology are shown with examples.
-S1 to S6 correspond to the locations of the six research areas in Table I: Wuhan, Hengyang, Shenzhen, Dallas, Kern, and Urumqi.
-TABLE I: Details of Six Research Areas
-  Satellite Task Location Date Range Area Size and Resolution
-  S1 Landsat-5 TM ETM+ SLC-off Wuhan, Hubei, China 2001-03 to 2003-10 (1720, 2040) , 30-m
-S2 Landsat-5 TM ETM+ SLC-off Hengyang, Hunan, China 2003-04 to 2005-05 (1200, 1200) , 30-m
-S3 Landsat-5 TM Thick Cloud Removal Shenzhen, Guangdong, China 2003-10 to 2004-11 (1680, 1680) , 30-m
-S4 Landsat-8 OLI Thick Cloud Removal Dallas, Texas, United States 2014-01 to 2014-11 (2040, 2040) , 30-m
-  S5 Landsat-7 ETM+ Real ETM+ SLC-off Kern, California, United States 2011-09 to 2011-11 (1720, 2040) , 30-m
-S6 Landsat-5 TM Real Thick Cloud Removal Urumqi, Xinjiang, China 2006-06 to 2009-08 (1440, 1440) , 30-m
-IV Experiment
-IV-A Experiment Settings
-To assess the performance of the proposed algorithm, we conducted experiments on both simulated and real-world datasets.
-Additionally, we investigated the effectiveness of each core component of the proposed method.
-IV-A1 Datasets
-Different datasets were tailored for different missing data reconstruction tasks.
-For instance, for simulated ETM+ SLC-off and thick cloud removal tasks, we selected two research areas each for model training and testing.
-For real-world SLC-off and thick cloud removal, we chose one research area each and evaluated the trained model within them.
-All research areas comprised multiple images captured at different times.
-We employed a sliding window approach with a stride of 60 and a size of 120$\times$120 to augment the sample amount, and then split them into training and testing sets.
-Specifically, Fig.
-6 illustrates the spatial distribution and examples of each research area, while Table I provides detailed information on their data sources, tasks, time ranges, etc.
-IV-A2 Comparison Algorithms and Evaluation Metrics
-For the ETM+ SLC-off and thick cloud cover problems, we compared our method with LLHM [21], WLR [24], and STS-CNN [37].
-These algorithms take a target image and a single temporal auxiliary image as input and synthesise a single reconstructed image, thus lacking the ability to handle long temporal auxiliary images, i.e., Dual-In Single-Out.
-In contrast, our proposed method leverages multi-temporal images with missing data as input and outputs reconstructed multi-temporal images in parallel, i.e., Multi-In Multi-Out.
-The simulated experiments utilized the mean peak signal-to-noise ratio (mPSNR) and the mean structural similarity index (mSSIM) for all spectral bands [6] as evaluation metrics.
-Furthermore, we repeated each experiment 5 times and used the average results as the final evaluation metric.
-IV-A3 Implement Detail
-For deep learning-based models, we utilized the Adam optimizer with parameters ($\beta_{1}=0.9$, $\beta_{2}=0.999$) for training, and set the batch size to 8.
+For each scale of MS2TAN, the expression of the loss function $\mathcal{L}^{\left(i\right)}$ corresponding to its intermediate output $\widetilde{Y}^{(i)}$ is shown in Eq.
+(25), where $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ denote the weights of the pixel-wise loss, structural loss and perceptual loss respectively.
+The multi-objective loss function $\mathcal{L}$ considers the intermediate results of each scale, using their mean as the overall loss, as shown in Eq.
+(26).
+$\mathcal{L}^{\left(i\right)}\left(\eta,y\right)={\lambda_{1}}\mathcal{L}_{\text{pixel-wise}}+{\lambda_{2}}\mathcal{L}_{\mathrm{struct}}+{\lambda_{3}}\mathcal{L}_{\mathrm{perceptual}}$ (25)
+$\mathcal{L}=\frac{1}{S}\sum_{i=1}^{S}\mathcal{L}^{\left(i\right)}\left(\widetilde{Y}^{(i)},Y\right)$ (26)
+IV Experiments
+IV-A Settings
+To assess the performance of the proposed model in different scenarios, we conducted quantitative and qualitative experiments on two datasets and compared it with some mainstream methods.
+In addition, we conduct ablation experiments on the key innovations and discuss the balance between effectiveness and efficiency for models with different sizes.
+IV-A1 Compared Algorithms
+Based on the type of inputs and outputs, common image restoration methods can be categorized into three main types: single-input-single-output (SISO), dual-input-single-output (DISO), and multi-input-multi-output (MIMO).
+The SISO algorithms are the simplest type, but they struggle to achieve excellent performance, so we do not compare this category with ours.
+The DISO algorithm takes the target image and a temporal reference image as input and synthesizes a single reconstructed image.
+For these methods, we use the less missing temporal neighbor of the target image as the reference.
+The DISO methods we compare include simple replacement (replace missing values by copying data from reference image), LLHM [21], WLR [24], and STS-CNN [37].
+The MIMO method utilizes a multi-temporal image containing missing data as input and outputs a reconstructed multi-temporal image, allowing the model to fully take into account temporal semantics.
+Compared methods of this type include last padding (Last), nearest neighbor padding (Nearest), and linear interpolation (Linear), which are commonly used in cartography [50], as well as the U-TILISE [40].
+IV-A2 Evaluation Metrics
+Through the sliding windows approach, we reconstructed and assessed data gaps across all time steps to facilitate a more comprehensive analysis.
+The quantitative experiments utilized mean absolute error (MAE), spectral angle mapper (SAM) [51], mean peak signal-to-noise ratio (PSNR), and mean structural similarity (SSIM) [47] as evaluation metrics.
+In addition, for real data gaps where ground truth is missing, we clearly show the comparison of restoration results for qualitative assessment.
+IV-A3 Implement Details
+For MS2TAN, we utilized the Adam optimizer with parameters ($\beta_{1}=0.9$, $\beta_{2}=0.999$) for training, and set the batch size to 8.
 The initial learning rate was set to $4\times 10^{-4}$, with a decay schedule of every 100 epochs, and an early stopping strategy of 30 epochs on the validation set to prevent overfitting.
+For the other models we compared, we conducted our experiments while following the original authors’ training settings and environment.
 The proposed MS2TAN was implemented with PyTorch 1.12 framework and trained with a NVIDIA GeForce RTX 3090 24GB GPU on a Ubuntu 20.04 environment.
-Additionally, the test code of MS2TAN is openly available on the Github at https://github.com/zzaiyan/MS2TAN.
-IV-B Simulated Experiment
-Figure 7: Experimental results of Simulated ETM+ SLC-off Restoration based on Research Area S1.
-(a) Ground truth (March 19, 2002).
-(b) Simulated ETM+ SLC-off.
-(c) LLHM+LPRM.
-(d) WLR+LPRM.
-(e) STS-CNN.
-(f) Ours.
-Figure 8: Experimental results of Simulated ETM+ SLC-off Restoration based on Research Area S2.
-(a) Ground truth (April 15, 2003).
-(b) Simulated ETM+ SLC-off.
-(c) LLHM+LPRM.
-(d) WLR+LPRM.
-(e) STS-CNN.
-(f) Ours.
-IV-B1 Simulated ETM+ SLC-off Restoration Experiment
-For the land satellite ETM+ SLC-Off problem, we conducted simulated experiments with synthesized missing data in research areas S1 and S2, as illustrated in Fig.
-7-8.
-In these figures, (a) and (b) represent the original image and the simulated SLC-off image, respectively, while (c)-(f) show the simulated recovery results using four methods—LLHM, WLR, STS-CNN, and the proposed MS2TAN.
-To provide a clearer view of the recovery results, we present magnified details next to each reconstruction.
-As shown in Fig.
-7-8 (c)-(e), all compared methods generate discontinuous fine features to some extent.
-This is because the temporal auxiliary images cannot fully cover the missing regions, thus requiring LLHM and WLR algorithms to rely on LPRM to fill the remaining gaps, resulting in highly blurred noise bands in the reconstructed results.
-Although the STS-CNN model based on an end-to-end strategy can repair these gaps, it still cannot accurately restore the original features.
-In contrast, the proposed MS2TAN, utilizing a Masked Spatial-Temporal Attention mechanism and integrating a “Pixel-Structure-Perception” Multi-Objective Optimization approach, achieves better restoration of texture and structural details, as shown in Fig.
-7-8 (f).
-Moreover, this method outperforms state-of-the-art methods in the quantitative evaluations presented in Table II, validating its strong capability in spatio-temporal integrated data utilization.
-Figure 9: Experimental results of Simulated Thick Cloud Removal based on Research Area S3. (a) Ground truth (October 17, 2003). (b) Simulated Thick Cloud. (c) LLHM. (d) WLR. (e) STS-CNN. (f) Ours.
-Figure 10: Experimental results of Simulated Thick Cloud Removal based on Research Area S4. (a) Ground truth (January 22, 2014). (b) Simulated Thick Cloud. (c) LLHM. (d) WLR. (e) STS-CNN. (f) Ours.
-TABLE II: Quantitative Evaluation Result of the Simulated SLC-off Restoration in Research Area S1 and S2
-  Dataset Area S1 Area S2
-Metric mPSNR($\uparrow$) mSSIM($\uparrow$) mPSNR($\uparrow$) mSSIM($\uparrow$)
-  LLHM [21] 32.424 0.8947 32.908 0.9049
-WLR [24] 34.374 0.9179 34.498 0.9214
-STS-CNN [37] 37.118 0.9446 37.761 0.9479
-$\textbf{MS}^{2}\textbf{TAN}$ (Ours) 38.255 0.9548 38.963 0.9555
-IV-B2 Simulated Thick Cloud Removal Experiment
-Similar to the simulated experiments of ETM+ SLC-off, we simulated thick cloud removal tasks in research areas S3 and S4, and reconstructed the data using the same four methods, resulting in the outcomes shown in Fig.
-9-10.
-Among them, LLHM exhibits severe texture discontinuities, while WLR produces numerous noise points.
-This is because these simple models struggle to handle highly complex nonlinear relationships between different times.
-In comparison, the STS-CNN model can more accurately reconstruct the texture and structural features of missing parts, but exhibits noticeable spectral transition phenomena at the edges of missing value regions.
-On the other hand, the proposed MS2TAN method introduces missing value hint tensors into the learning process, enhancing the model’s capability to handle irregularly distributed missing values.
-Ultimately, it performs well in maintaining consistency in texture, structure, and spectral properties, achieving the best metrics in the quantitative evaluation presented in Table III.
-TABLE III: Quantitative Evaluation Result of the Simulated Thick Cloud Removal in Research Area S3 and S4
-  Dataset Area S3 Area S4
-Metric mPSNR($\uparrow$) mSSIM($\uparrow$) mPSNR($\uparrow$) mSSIM($\uparrow$)
-  LLHM [21] 34.775 0.9439 34.714 0.9370
-WLR [24] 39.003 0.9639 38.261 0.9614
-STS-CNN [37] 39.515 0.9711 38.753 0.9662
-$\textbf{MS}^{2}\textbf{TAN}$ (Ours) 39.802 0.9753 39.272 0.9693
-Figure 11: Experimental results of Real ETM+ SLC-off Restoration based on Research Area S5.
-(a) SLC-off Image 1 (October 23, 2011).
-(b) SLC-off Image 2 (November 8, 2011).
-(c) LLHM+LPRM.
-(d) WLR+LPRM.
-(e) STS-CNN.
-(f) Ours.
-Figure 12: Experimental results of Real Thick Cloud Removal based on Research Area S6.
-(a) Image with clouds (August 5, 2007).
-(b) Temporal image without clouds (July 20, 2007).
-(c) LLHM.
-(d) WLR.
-(e) STS-CNN.
-(f) Ours.
-IV-C Real Data Experiment
-IV-C1 SLC-off Restoration Experiment
-The experimental results of real SLC-off images in research area S5 are shown in Fig.
+Additionally, the test code of MS2TAN is openly available on the Github at https://github.com/CUG-BEODL/MS2TAN.
+IV-B Datasets
+We conducted experiments on two datasets, Landsat and EarthNet2021 [52]. Their details are described in Table VI.
+TABLE I: Quantitative Evaluation Results under Dual-Temporal Inputs Compared to the DISO Methods, using the Landsat and EarthNet2021 datasets.
+The results for the best-performing method for each metric are highlighted in bold.
+  Method Type Landsat EarthNet2021
+$\downarrow$ MAE $\downarrow$ SAM $\uparrow$ PSNR $\uparrow$ SSIM $\downarrow$ MAE $\downarrow$ SAM $\uparrow$ PSNR $\uparrow$ SSIM
+  Replace DISO 0.0210 2.81 32.67 0.8913 0.0146 3.10 34.04 0.9452
+LLHM [21] 0.0161 1.95 34.64 0.9104 0.0121 2.59 35.54 0.9566
+WLR [24] 0.0110 1.33 36.88 0.9312 0.0113 2.41 36.08 0.9605
+STS-CNN [37] 0.0091 1.12 37.79 0.9423 0.0102 2.18 37.07 0.9647
+  MS2TAN (Ours) DIDO 0.0081 1.02 38.54 0.9503 0.0095 2.04 37.60 0.9671
+Figure 5: Visual comparison results between MS2TAN and the DISO algorithms on the Landsat dataset, with dead pixels shown in black.
+Row 2 details the highlighted area in row 1.
+From left to right: (a) Input 1 (target).
+(b) Input 2 (reference).
+(c)-(g) restoration results of each method.
+(h) Ground truth.
+IV-B1 Landsat
+The Landsat dataset contains a large number of Top-of-Atmosphere (TOA) time-series captured by Landsat-5/7/8 satellites, with approximately uniform time intervals.
+The original images are cropped into small patches of size $120\times 120$ pixels with a spatial resolution of 30 meters.
+Each image includes 6 bands (Blue, Green, Red, NIR, SWIR1, and SWIR2), as well as the SLC-off or cloud mask provided in the QA band.
+In terms of spatial extent, we focus on sequences from multiple study regions in China and the United States.
+IV-B2 EarthNet2021
+The EarthNet2021 [52] dataset comprises over 28,000 Sentinel-2 TOA time-series observations from the European continent, with uniformly spaced time intervals.
+Each image consists of 4 bands (Blue, Green, Red and NIR), at a size of $128\times 128$ pixels, downsampled to a spatial resolution of 20-meters, accompanied by cloud probability maps.
+We utilized 80% of sequences from the training split for training purposes, with the remaining set aside for validation.
+For testing, we used sequences from the iid test split.
+IV-B3 Preprocessing
+We follow the preprocessing protocol of EarthNet2021 to crop the values of the optical images to the range $\left[0,10~{}000\right]$ and then normalize them to the unit range $\left[0,1\right]$.
+To match the multi-scale patch division strategy of MS2TAN, we resize the input images to a common multiple of the patch size of the MSTA at different scales.
+In our experiments, we use three scales (patch size is 8, 10, and 12 respectively) which have a least common multiple of 120.
+We therefore centrally cropped the images in the EarthNet2021 dataset to $120\times 120$ pixels, in line with the Landsat dataset.
+IV-C Results
+TABLE II: Quantitative Evaluation Results under Multit-Temporal Inputs Compared to the MIMO Methods, using the Landsat and EarthNet2021 datasets.
+The results for the best-performing method for each metric are highlighted in bold.
+  Method Type Landsat EarthNet2021
+$\downarrow$ MAE $\downarrow$ SAM $\uparrow$ PSNR $\uparrow$ SSIM $\downarrow$ MAE $\downarrow$ SAM $\uparrow$ PSNR $\uparrow$ SSIM
+  Last MIMO 0.0223 3.14 32.11 0.8862 0.0148 3.17 33.68 0.9439
+Nearest 0.0198 2.54 33.28 0.8957 0.0128 2.74 35.04 0.9534
+Linear 0.0103 1.26 37.20 0.9348 0.0110 2.35 36.48 0.9620
+U-TILISE [40] 0.0082 1.05 38.45 0.9491 0.0086 1.87 38.29 0.9702
+MS2TAN (Ours) 0.0074 0.96 39.01 0.9552 0.0078 1.71 38.91 0.9728
+Figure 6: Visual comparison results between MS2TAN and the MIMO algorithms on the EarthNet2021 dataset, with thick cloud shown in white.
+Each row shows a time-ordered image sequence.
+From top to bottom: input sequence (row 1), restoration results of each method (rows 2-6), and ground truth (row 7).
+To quantitatively compare restoration results, we evaluated our proposed approach as well as mainstream methods on synthetic data gaps.
+Specifically, we performed comparison experiments under two different conditions: dual-temporal inputs (vs.
+the DISO algorithms), and multi-temporal inputs (vs.
+the MIMO algorithms).
+The quantitative evaluation results on the Landsat and EarthNet2021 datasets are shown in Table I and II, and visual eaxamples are illustrated in Figs.
+5 and 6.
+Furthermore, Figs.
+7 and 8 shows the restoration results in real data gaps to validate the generalisation performance of MS2TAN.
+IV-C1 Comparison under Dual-Temporal Inputs
+We conducted comparison experiments with the DISO algorithms using synthetic data gaps in the Landsat and EarthNet2021 datasets.
+As illustrated in Fig.
+5, (a) and (b) represent the two images with data gaps used as input, while (c)-(g) sequentially display the recovery results obtained from five methods, with (h) showing the ground truth of the target image.
+As depicted in (c)-(f), all compared methods to some extent produced discontinuous fine features.
+This is because the time-assisted images were unable to completely cover the missing regions, leading the LLHM, and WLR to rely on LRPM (Laplacian Prior Regularization Method [24]) to fill in the remaining gaps, resulting in very blurry noise bands in the reconstruction results.
+Although the end-to-end strategy-based STS-CNN model can repair these blanks, it fails to accurately restore the original features.
+In contrast, the proposed MS2TAN, integrating a “Pixel-Structure-Perception” Multi-Objective Joint Optimization approach, achieves superior texture and structural detail recovery, as shown in (g).
+In the quantitative evaluation shown in Table I, MS2TAN outperforms other comparative methods across all evaluation metrics.
+Compared to the previous SOTA method STS-CNN, MS2TAN exhibits a 10.99%/6.86% (in two datasets respectively) reduction in MAE and a 0.75 dB/0.53 dB increase in PSNR, showcasing its robust capability in utilizing spatiotemporal data.
+IV-C2 Comparison under Multi-Temporal Inputs
+Figure 7: Visual evaluation of restoration results for real ETM+ SLC-off images from the Landsat dataset, with dead pixels shown in black.
+The first and second rows show the observed sequence and the restoration results of the MS2TAN, respectively.
+Figure 8: Visual evaluation of restoration results for real Thick Cloud Cover images in the EarthNet2021 dataset, with clouds shown in white.
+The first and second rows show the observed sequence and the restoration results of the MS2TAN, respectively.
+Similar to the dual-temporal input experiment, we conducted comparative experiments with the MIMO algorithms under multi-temporal input conditions.
+Visual comparative results are illustrated in Fig.
+6, displaying, from top to bottom: the input image sequence, the restoration outcomes of five methods, and the ground truth.
+Since traditional linear methods rely on time-series information, they cannot restore regions in the input sequence where all time steps are missing.
+In contrast, U-TILISE can recover data gaps of any shape, but still exhibits significant color shifts in irregularly shaped gaps, reducing the quality of the results, as shown in the first three columns of Fig.
+6.
+Benefiting from the Masked Spatial-Temporal Attention mechanism, MS2TAN can effectively handle the edges of missing regions, achieving more harmonious seams.
+Notably, the Multi-Objective Optimization makes MS2TAN perform better in restoring structural and textural details in the missing regions.
+In some extreme cases (e.g., the upper-left region of the first 8 columns), due to a lack of sufficient information, MS2TAN may not accurately reproduce the actual changes but can still output coherent images with natural color transitions.
+Compared to the U-TILISE, MS2TAN exhibits a 9.76%/9.30% reduction in MAE, a 0.56 dB/0.62 dB increase in PSNR, and achieves higher visual consistency.
+IV-C3 Restoration of the Real Data Gaps
+To evaluate the generalization performance of the proposed MS2TAN, we applied pre-trained model weights to real ETM+ SLC-off sequences and thick cloud-covered sequences.
+Due to the lack of ground truth reflectance data, the assessment was conducted based solely on visual inspection.
+Visualization of the restoration results is provided in Figs.
+7 and 8.
+In the restored ETM+ SLC-off images, the roads and rivers, which are divided by striped missing areas, are rationally connected in the restoration results with the original structure maintained.
+Obvious ground surface changes brought by long time intervals are also taken into account.
+MS2TAN utilizes Masked Spatial-Temporal Attention to learn the distribution of missing data, thereby capturing temporal evolution patterns from neighboring frames and fitting spatial textures, achieving high-precision data restoration.
+For the thick cloud-covered images, the method accurately restored all missing data, including both large-scale and fragmented gaps.
+Based on the assistance of long time-series, continuous large-area data gaps can also be repaired reasonably well and keep the original evolutionary trend.
+Irregularly shaped data gaps (e.g., the images in the sixth column in Fig.
+8) are repaired to blend in with the spatial domain, preserving high textural and structural coherence.
+By applying Masked Spatial-Temporal Attention and Multi-Objective Joint Optimization, MS2TAN exhibits excellent performance in handling complex and large-scale data gaps, which often present challenges in traditional data recovery methods, as evident in the detailed visual evaluation.
+IV-D Validation Studies
+To validate the effectiveness of the current architecture, we experimented with some variants of MS2TAN and obtained quantitative results as shown in Table III.
+We discuss each of the mechanisms used below.
+TABLE III: Quantitative Evaluation of Different MS2TAN Variants that derived from whether Separated Spatial-Temporal Attention (Sep.
+Attn), Attention Mask Mechanism (Attn Mask), and Multi-scale Restoration Strategy (Multi-Scale) are enabled.
+Statistics information and evaluation metrics are provided on the right.
+All experiments are performed under Multi-Temporal Inputs.
+       Features Statistics Landsat EarthNet2021
+Sep. Attn Attn Mask Multi-Scale Params (M) FLOPs (G) $\downarrow$ MAE $\downarrow$ SAM $\uparrow$ PSNR $\uparrow$ SSIM $\downarrow$ MAE $\downarrow$ SAM $\uparrow$ PSNR $\uparrow$ SSIM
+  4.33 8.756 0.0074 0.96 39.01 0.9552 0.0078 1.71 38.91 0.9728
+  4.33 14.626 0.0073 0.95 39.08 0.9560 0.0083 1.81 38.53 0.9714
+4.33 8.756 0.0096 1.18 37.47 0.9384 0.0112 2.38 36.37 0.9609
+6.61 10.130 0.0086 1.07 38.09 0.9459 0.0095 2.05 37.60 0.9672
+6.61 16.473 0.0101 1.25 37.24 0.9356 0.0127 2.73 35.19 0.9538
+IV-D1 Separated Spatial-Temporal Attention
+The original ViT, after embedding images into token sequences, feeds the entire sequence into a multi-head self-attention module.
+However, the computational complexity of self-attention is proportional to the square of the sequence length.
+For a sequence of remote sensing images ($t$ images, $n$ patches per image), the computational complexity is as high as $O\left(t^{2}\times n^{2}\right)$.
+Whereas, separating temporal attention from spatial attention and computing them sequentially reduces the complexity to $O\left(t^{2}+n^{2}\right)$.
+Although the perceptual field of the separated spatial-temporal attention is reduced in a single operation, we alternate multiple temporal and spatial attention modules to spread the perceptual field to the whole spatiotemporal range.
+Meanwhile, the extension of the perceptual field along the temporal axis and the spatial plane matches the characteristics of the time-series remote sensing images, which have small displacement in time and large spatial correlation, and can fully integrate the spatial and temporal information.
+As shown in the first two rows of Table III, the use of separated spatial-temporal attention reduces the computational effort by 40.1% with almost the same accuracy in the Landsat dataset, while the accuracy in EarthNet2021 is slightly improved.
+Figure 9: Visualization of Temporal Attention Distribution in the first MSTA unit.
+We present the attention scores for repairing the 3rd frame (highlighted with a red box in row 1) of the sequence.
+It is observed that, after incorporating the mask, the temporal attention is no longer restricted to the vicinity of the target time but instead focuses on a broader range of valid pixels.
+From top to bottom: Input (row 1), w/o Mask (rows 2-5), and w/ Mask (rows 6-9).
+Figure 10: Visualisation of Spatial Attention Distribution in the first MSTA unit.
+We show the average attention scores for imputing each patch.
+With the mask, spatial attention focuses on the edges of the valid pixels, extracting textural and structural features, rather than spreading around as it would be without the mask.
+From top to bottom: Input (row 1), w/o Mask (rows 2-3), and w/ Mask (rows 4-5).
+Figure 11: Ablation Experiment Results for Optimal $C_{\text{max}}$ Value in MSTA.
+The horizontal axis represents different $C_{\text{max}}$ values, and the vertical axis represents the SSIM in the EarthNet2021 dataset.
+IV-D2 Attention Mask in MSTA
+To verify the effectiveness of the attention mask in MSTA, we visualized the distribution of attention.
+Fig.
+9 shows the distribution of temporal attention with and without the attention mask, while Fig.
+10 illustrates the spatial attention distribution.
+Without the mask, certain attention heads focus excessively on patches with missing data, while others focus excessively on their own patch, providing no additional information to impute missing values and leading to unnecessary computational costs.
+With the mask mechanism in place, the attention heads no longer focus on missing-data patches or themselves, but instead extract information from spatiotemporal domains, obtaining a larger effective receptive field to achieve higher Imputation accuracy.
+Within the MSTA framework, the maximum allowable missing rate, $C_{\text{max}}$, is a crucial hyperparameter.
+Setting it too high can cause patches with missing data to distort token distribution, while a too-low value may result in information loss.
+Consequently, we conducted ablation experiments on different values of $C_{\text{max}}$, with the results shown in Fig.
 11.
-In this figure, (a) and (b) depict two ETM+ SLC-off images observed on October 23, 2011, and November 8, 2011, respectively, while (c)-(h) display the outputs of four algorithms.
-Similar to the simulated SLC-off reconstruction experiment, single temporal auxiliary images cannot fully cover the gaps.
-Thus, after LLHM and WLR processing, the remaining gaps are filled using the LPRM algorithm.
-STS-CNN can directly repair the entire image but exhibits discontinuous texture features similar to the simulated experiment.
-Thanks to the “Pixel-Structure-Perception” Multi-Objective Joint Optimization method, the proposed MS2TAN method preserves more structural details and produces more consistent reconstruction results.
-IV-C2 Thick Cloud Removal Experiment
-The experimental results of thick cloud coverage reconstruction in region S6 are shown in Fig.
+The findings indicate that an optimal $C_{\text{max}}$ yields the best reconstruction accuracy, and thus $C_{\text{max}}$ is set to 0.5 for subsequent experiments in this paper.
+IV-D3 Multi-scale Restoration
+Figure 12: Visualisation of intermediate results for multi-scale restoration.
+The values above indicate the MAE compared to the ground truth.
+(a) Input.
+(b)-(d) Intermediate outputs at each scale.
+(e) Ground truth.
+MS2TAN employs a multi-scale restoration strategy to refine the restoration results layer by layer.
+To verify its effectiveness, we take the SLC-off reconstruction task as an example and compare the intermediate outputs of each layer in Fig.
 12.
-In Fig.
-12 (a), an image containing thick clouds is presented, while (b) depicts the temporal auxiliary image.
-Four algorithms identical to those used in the SLC-off reconstruction experiment were employed to reconstruct the images, yielding results displayed in (c)-(f).
-It can be observed in Fig.
-12 (a)-(b) that for extensive cloud coverage and cloud shadows, the results of LLHM and WLR models suffer from significant loss of texture and structural features, presenting as blurry patches.
-This is attributed to the inconsistency between the reconstructed texture details in the cloud region and the surrounding non-cloud areas, which the aforementioned models fail to fit due to their incapability to capture the nonlinear relationship between different temporal data.
-In comparison, the STS-CNN method exhibits better performance in texture restoration, yet still shows noticeable spectral discontinuities at the boundaries of missing regions.
-On the other hand, the proposed MS2TAN method fully preserves texture and structural details with less spectral distortion compared to STS-CNN, resulting in high-quality thick cloud removal images.
-IV-D Validation Study
-IV-D1 Masked Spatial-Temporal Attention
-To validate the effectiveness of the proposed Masked Spatial-Temporal Attention (MSTA) mechanism, we compared models with and without the adoption of this mechanism, and plotted their mSSIM iteration curves on the validation set, as shown in Fig.
-13.
-It can be observed that with the assistance of the attention mask mechanism, the model exhibits faster convergence speed and higher fitting accuracy.
-This confirms that the attention mask mechanism effectively enhances the model’s feature extraction capability and training speed.
-Figure 13: Ablation experiments for the proposed MSTA mechanism.
-The blue and red curves respectively depict the variations of mSSIM with training rounds for models with and without the adoption of the attention mask mechanism in Area S1.
-IV-D2 Multi-scale Restoration
-The MS2TAN employs a multi-scale restoration strategy to process input images.
-To validate its effectiveness, we take the SLC-off reconstruction task as an example and compare the intermediate outputs of each layer.
-The results are shown in Fig.
-14.
-In the figure, the missing regions of the input image are filled with zeros.
-After Coarse Embedding, MFE, and Unembedding, a coarse repair result is obtained.
-This result reconstructs most of the texture and structural features and balances the significant distribution differences between the missing regions and other areas in the input.
-However, there are still spectral discontinuities at the edges of the missing regions.
-After several layers of restoration, the final refined repair result is obtained.
-This result refines the previous intermediate results, with clearer textures, consistent spectra, and smoother integration of the repair area boundary with other areas.
-Figure 14: Ablation experiments for multi-scale restoration.
-(a)-(c) represent observations from three different times, processed in parallel by MS2TAN after composing the time-series remote sensing images.
-From top to bottom, they represent: input with missing values, coarse result, and fine result.
-IV-D3 Gains from long time-series image inputs
-A highlight of the proposed model is its capability to utilize longer time-series images for assistance in restoration, leading to more accurate reconstruction results.
-We designed experiments to compare different lengths of time-series inputs to quantitatively evaluate the accuracy gains from longer time-series inputs.
-The experimental results are shown in Fig.
-15.
-From Fig.
-15, it can be observed that as the length of the input time series increases from 2, there is a significant improvement in mPSNR and mSSIM metrics.
-However, when the time series length reaches 5 or longer, the change in accuracy becomes less significant.
-Therefore, longer time-series inputs can significantly improve the accuracy of missing value reconstruction, but there is a certain limit.
-Hence, we can find a balance point between reconstruction accuracy and algorithm flexibility.
-Figure 15: Analysis of the input time-series length.
-The left and right tables record the variations of mPSNR and mSSIM with the length of the time-series, respectively, and distinguish the results of four research areas with different colors.
+With the gradual deepening of the restoration stages, the restoration results are gradually refined to better capture the details and structural features in the image.
+Specifically, the coarse repair stage addresses global textures and structural issues, while the fine repair stage further optimizes local details and spectral consistency.
+However, despite the effectiveness of the multi-scale approach, introducing too many scale levels leads to diminishing returns, with further improvements becoming negligible or even introducing overfitting and increased computational costs.
+Therefore, selecting an appropriate number of scale levels is crucial.
+IV-D4 Multi-Objective Joint Optimization
+TABLE IV: Ablation Analysis for Multi-Objective Joint Optimization, where $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ denote the weights of the pixel-wise loss, structural loss and perceptual loss respectively.
+         Loss weights       EarthNet2021
+  $\lambda_{1}$ $\lambda_{2}$ $\lambda_{3}$downarrow MAE $\downarrow$ SAM $\uparrow$ PSNR $\uparrow$ SSIM
+  1 - - 0.0120 2.55 35.76 0.9572
+0.9 0.1 - 0.0091 1.97 37.94 0.9686
+0.9 - 0.1 0.0102 2.20 37.09 0.9645
+- 0.5 0.5 0.0210 4.43 31.72 0.9241
+  0.9 0.05 0.05 0.0078 1.71 38.91 0.9728
+In Multi-Objective Joint Optimization, the weights assigned to each loss function are critical.
+To investigate this, we conducted an ablation study on the allocation of loss weights.
+Table IV presents different weight configurations for pixel-wise loss ($\lambda_{1}$), structural loss ($\lambda_{2}$), and perceptual loss ($\lambda_{3}$), along with their corresponding results on the EarthNet2021 dataset.
+The findings show that combining pixel-wise loss with either structural or perceptual loss leads to significant performance improvements (cf.
+rows 2-3), with the best results achieved when both are combined.
+However, when only structural and perceptual losses are applied, the model struggles to converge, resulting in poor performance (cf.
+row 4).
+Therefore, pixel-wise loss plays a dominant role in multi-objective optimization, and its combination with structural and perceptual losses is essential for achieving high restoration accuracy and visual consistency.
+TABLE V: Hyperparameter settings and Evaluation of MS2TAN with different sizes.
+The hyperparameters include the restoration scales number ($S$), patch size ($P$), embedding dimensions ($d_{\text{emb}}$), attention heads number ($h$), dimensions of vectors Q, K and V in self-attention ($d_{\text{qkv}}$), and the MSTA units number ($L$), where bracketed lists indicate the hyperparameters for each scale.
+  Model Hyperparameters Statistics EarthNet2021
+$S$ $P$ $d_{\text{emb}}$ $h$ $d_{\text{qkv}}$ $L$ Params (M) FLOPs (G) $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM
+  MS2TAN 3 (12, 10, 8) (256, 192, 128) (8, 6, 4) (32, 32, 32) (2, 2, 2) 4.33 8.756 0.0078 1.71 38.91 0.9728
+  MS2TAN-L 3 (12, 10, 8) (384, 256, 192) (8, 6, 4) (48, 48, 48) (4, 4, 4) 17.04 34.363 0.0076 1.67 39.03 0.9733
+MS2TAN-S 2 (12, 10) (192, 128) (8, 6) (24, 24) (2, 2) 2.07 3.567 0.0089 1.92 38.07 0.9692
+IV-D5 Balance between Effectiveness and Efficiency
+Scalability is a key strength of ViT-based models, and MS2TAN is no exception.
+To strike a balance between effectiveness and efficiency, we adjusted several critical hyperparameters to obtain two MS2TAN variants and evaluated them on EarthNet2021 dataset (see Table V).
+The experimental results reveal that when the model is too small (MS2TAN-S), performance significantly degrades; conversely, when the model is too large (MS2TAN-L), there is only a slight performance gain, while the computational cost increases drastically, making the approach less efficient.
+Overall, the current configuration of parameters and computational load represents an optimal “sweet spot” where good performance is maintained without excessive computational overhead.
 V Conclusion
-In this paper, we proposed a novel method for time-series remote sensing image reconstruction based on a Multi-scale Masked Spatial-Temporal Attention Network.
-The proposed method reconstructs various missing value scenarios, including the Landsat ETM+ SLC-off problem and removal of thick cloud coverage in remote sensing images.
-Compared to existing methods, the proposed method improves the efficiency of spatio-temporal information utilization, enabling the mutual assistance of multiple degraded time-series images for reconstruction and achieving higher restoration accuracy through more auxiliary information.
-Additionally, the application of the MSTA mechanism and Multi-Objective Joint Optimization further enhances the texture and structural consistency of the reconstruction results.
-Experimental results demonstrate significant improvements in visual effects and restoration accuracy compared to mainstream methods.
-Although the proposed method performs well in reconstructing the ETM+ SLC-off problem and thick cloud removal, there are still some inevitable limitations.
-For example, the model needs to simulate various shapes of cloud-like missing regions in the dataset and learn the characteristics of missing value distribution extraction sufficiently to achieve better cloud removal effects, which requires more training time for the model.
-In our future research, we can incorporate network components pre-trained on large-scale datasets into the model to improve its convergence speed.
+This paper presents a novel approach to reconstruct time-series remote sensing images using the Multi-Scale Masked Spatial-Temporal Attention Network (MS2TAN).
+MS2TAN takes an image sequence with arbitrary data gaps as input and generates a clear and complete time-series.
+Compared with the existing methods, the proposed Masked Spatial-Temporal Attention (MSTA) mechanism improves the efficiency of spatiotemporal information utilization, reduces the high computational cost of the original ViT, and obtains higher restoration accuracy.
+In addition, the application of the Multi-scale Restoration strategy and Multi-Objective Joint Optimization further enhances the texture and structural consistency of the reconstructed images.
+MS2TAN achieves higher imputation accuracy than the mainstream methods in quantitative experiments and exhibits better visual effects in eliminating real data gaps.
+Ablation experiments also confirm the contributions of primary innovations to the results.
+Despite the effectiveness of the proposed method in addressing ETM+ SLC-off and thick cloud cover, there are some unavoidable limitations.
+For instance, the model needs to simulate various shapes of data gaps in the dataset to adequately learn to extract the features of the missing value distribution in order to achieve better restoration results, which leads to the model needing more time for training.
+Therefore, future research could incorporate network components pre-trained on large-scale datasets to improve model convergence speed and generalization performance.
 Acknowledgments
-This work was supported in part by the National Key R&D Program of China (Project No. 2022YFC3800700); in part by the National Natural Science Foundation of China (No. U21A2013).
+This work was supported in part by the National Natural Science Foundation of China (No. 42471505); in part by the National Key R&D Program of China (Project No. 2022YFC3800700).
+Table VI details the statistics of the dataset we used, including key information such as data source, resolution, and number of sequences.
+TABLE VI: Details of the Landsat and EarthNet2021 Datasets.
+  Landsat       EarthNet2021
+        Data source Landsat-5/7/8       Sentinel-2
+Restoration task SLC-off & Thick Cloud Thick Cloud
+Spatial scope China & USA Europe
+Temporal scope 2001 - 2011 2016 - 2020
+        Spectral bands 6 bands       4 bands
+Image size 120 $\times$ 120 128 $\times$ 128
+Spatial Resolution 30-m 20-m
+Temporal length 10 30
+Temporal interval average 30 days constant 5 days
+        Number of samples 11913       28123
+– Train 7625 18955
+– Valid 1905 4949
+– Test 2383 4219